A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2025–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

06 January 2025 Version 1.06

Title word cross-reference

(2 + 1) [MO25]. 0 [Kar25b]. 1 [AKKM25, Kar25b]. 2 [LMZ25]. AXB = C [NZ25]. B [GP25]. c [XQW25]. C^{0} [GS25]. C^{1} [GP25]. H(curl^{2}) [WLZZ25]. H_{∞} [SC25]. k [DJ25, GJ25, JFP25]. k/n(G) [YSP25]. L1 [CG25]. L^{2} [ZC25a, Xie25]. l_{1/2} [CLJX25]. L_{2} [LSPV25]. L_{∞} [ZZJJ25]. R [CdlHG25]. M [NDSZ25]. n [HY25b, Ran25]. n: G [DJ25]. R [XNL25]. R^{n} [BS25a, CA25b]. R_{0} [SWD25]. ρ [TST25]. W̃_{2}(m, m − 1) [SHK25]. φ [TST25]. w [AAN25].

-conforming [WLZZ25]. -dimensional [MO25, Ran25]. -means [XQW25]. -norm [Xie25]. -out-of- [DJ25]. -out-of-n [GJ25]. -person [HY25b]. -record [JFP25].

-splines [GP25]. -tensors [NDSZ25]. -type [SWD25].

1 [Ano25a, Ano25b, Ano25c]. 15

[Ano25d, Ano25e, Ano25f].

2025 [Ano25g, Ano25q, Ano25a, Ano25d, Ano25r, Ano25b, Ano25e, Ano25f, Ano25c].

421 [MR25].

absolute [LC25a]. accelerated [HHF25, ZLZ25]. accuracy [BAPPF25, MCL25]. accurate [YWW25]. actuarial [HY25a]. ADAM [SWWY25]. ADAM-type [SWWY25]. Adaptive [MWTW25, JCCZ25, JWS^{+}25, OAO25, TGDC25, ZMG25, ZLC^{+}25]. adaptivity [BS25b, KS25b, YTT25]. additive [Ran25]. advection [FBM25, XGYH25]. adversarial [AGGM25]. affinely [Sah25]. against [SBAM25]. agent [LSCZ25]. algorithm [AK25a, CLJX25, CLY25b, CLQY25, CXLP25, FA25, GLTW25, HBG25, IRS25, JJM25, LZFP25, LC25a, LJY25, LY25, LR25, LZZ^{+}25b, OAO25, SQLZ25, SWWY25, SWD25, TZ25b, WKQ25, XNL25, XLLJ25, ZLZ25, ZLC^{+}25, ZHLD25]. algorithms [FBZ25, HHF25, JWS^{+}25, LSCZ25, MG25, NZ25, SSLP25, ZS25]. Allen [HKH^{+}25, XC25, ZSYZ25]. Almost [CCLH25, BS25c]. Alpha [PW25]. Alpha-robust [PW25]. alternating [XNL25, NDSZ25]. American [ZZ25]. Analysis [DLR25, JFP25, LCZ25, WZTL25, ADY25, BFQ25, CG25, FADA25, Kar25a, LW25, LTZ25, LP25b, LGDN25, LMZ25, LR25, SX25, SDS25, SGRS25, SBAM25, TC25, WCW25, YSP25, ZLS25, ZSYZ25, ZC25b, ZFCM25]. analytic [SH25]. analyzing [HLL25]. Anderson [NDSZ25]. anisotropic [WX25]. annular [YWW25]. Appl [MR25]. applicability [LP25b]. Application [CA25b, FKVZ25, BRPP25, CLY25b, CLQY25, CXLP25, FBM25, GP25, IRS25, JN25, LC25a, LY25, MZL25, TSM25, TA25, XQW25, Yao25]. applications [HY25a, SWD25, ZHLD25]. applied [MR23, MR25, XGYH25]. approach [ADJL25, GBB25, HBG25, SDS25, SGRS25]. Approaches [ZFCM25]. Approximate [KSC25, ZNL25]. approximates [Xie25]. approximating [SHK25]. Approximation [KCM25, AKKM25, CA25a, CdlHG25, HJ25, JN25, Öza25, SZD25, TC25, XNL25, XLLJ25, YC25]. approximations [GHZ25]. April [Ano25g]. arbitrary [AM25, LC25b, PY25]. argument [FA25, Pho25]. argument-based [FA25]. arising [COPPS25]. arrangement [MT25]. arrival [BC25]. Asian [AAN25]. assets [LG25, ZFCM25]. assimilation [DLR25, WJ25]. associated [BCD25]. asymptotic [ADJL25, FLS25, ZC25b]. augmented [HHF25, ZNL25]. authentic [LGDN25]. autoregressive [HY25a]. average [XLLJ25].

B [LWL^{+}25, RK25]. B-spline [LWL^{+}25, RK25]. Backward [KKV25, LTZ25, CP25, IRS25, YY25]. Banach [OAO25]. bank [ZFCM25]. barrier [HKY25]. based [ADJL25, AS25, CKC25, CLY25a, CLY25b, FA25, FWL25, GLTW25, JFP25, JJM25, JWS^{+}25, Kar25a, KS25c, LP25a, LZFP25, LWL^{+}25, LRG25, LG25, LJY25, LWZZ25, MS25, MWTW25, MHO25, MG25, OAO25, RK25, WCS25, XJQ25, XNL25, You25, YPLH25, ZLC^{+}25]. Basis [GCN25, KCM25, MHO25, WYZ25]. batch [SSLP25]. Bayesian [SDS25, YXM25]. BDF2 [YTT25]. behavior [KKV25]. behavioral [COPPS25]. Benjamin [KM25]. Bessel [LC25b, ZC25b]. between [Xie25]. Bézier [AM25]. Bézier-Smart [AM25]. bi [LG25]. bi-objective [LG25]. bicriteria [SSLP25]. bidiagonalization [LTZ25]. bifurcation [Kar25b, MO25]. biharmonic [BCD25]. bilinear [SZD25, XJQ25]. BINO [YXM25]. bioconvective [UX25]. biological [FBM25]. Biot [KL25]. biperiodic [DdF25]. biquaternion [CKC25]. bivariate [TA25]. Block [GH25, AK25b, BKST25, BS25c, CLJX25, KSC25, NZ25, SHL^{+}25, TGDC25]. block-by-block [KSC25]. blurred [JJM25]. Board [Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p]. Bona [KM25]. book [DG25]. bordered [BS25c]. bound [SWD25]. boundary [BCD25, LFH25, LP25b, XC25]. bounded [FKVZ25, HY25b]. Bounds [CP25, BRPP25, SH25]. Bregman [LY25]. Brinkman [JLC25, LZZ25a]. Brownian [AAN25, ZZX25]. Broyden [TZ25b]. Broyden-like [TZ25b]. bubble [GS25]. bubble-enriched [GS25].

Cahn [HKH^{+}25, XC25, ZSYZ25]. calculus [HuR25, Öza25, Sak25]. Can [NT25]. candidate [LZFP25]. capabilities [BDvdHV25]. cardiac [RNC25]. cardinality [WKQ25]. Carlo [SDS25]. case [Hag25]. Cauchy [LF25]. causality [BDvdHV25]. causality-preservation [BDvdHV25]. chain [HBG25]. chance [CLY25a]. chaos [Kar25b, MO25]. chaotic [ADY25]. Charged [Yin25]. Charged-Particle [Yin25]. Chebyshev [Moi25, PM25b]. chevron [KKV25]. class [Buo25, TSM25, WX25, WCS25, XHLH25]. classical [JN25]. clustering [XQW25]. coefficients [HHF25, YY25]. cold [MHWY25]. collected [DJ25]. Collocation [COPPS25, AD25, GHD25, HRST25, NKN25]. Color [CKC25]. combined [Che25, HY25a]. common [CXLP25]. communication [LSCZ25]. commutative [Hag25, YY25]. compact [ZZJJ25]. Comparison [Sak25, HRO^{+}25]. competing [WZTL25, YSP25]. complementarity [CLY25b, FWL25, HV25, MG25, WCS25, XLLJ25]. complementary [WZTL25]. completion [AS25]. complex [FKVZ25, FA25, LCZ25, MG25]. complex-valued [MG25]. Complexity [ZS25]. components [DJ25, GJ25]. composite [LGDN25]. compressed [CLJX25]. compression [SHL^{+}25]. compressive [ZHLD25]. Comput [MR25]. computation [CLZ25, GR25, Pho25, Sak25]. Computational [UX25, AD25]. computations [ZWSC25]. computed [LSPV25]. Computing [JVW25, SHL^{+}25, Yao25]. condition [XC25]. conditions [LP25b, Ran25, Ton25, Xie25, ZHLD25]. conduction [OTD^{+}25]. conductivity [Sai25]. cone [SWD25]. conformable [GLTW25, MO25]. conforming [WLZZ25]. conjugate [AMRS25, JWS^{+}25, ZHLD25]. connected [CS25]. connectivity [ZLY25]. conservation [ADJL25, FC25]. conservative [Yin25]. conserved [ZC25a]. conserving [ZMG25]. consistency [MCL25]. consistent [TGDC25]. constant [TA25]. constants [LZZ25a]. constrained [CLJX25, CLY25a, GS25, HHF25, JJM25, LS25, WKQ25]. constraint [GHZ25, LYC25, ZS25]. constraints [LG25, RNC25, ZLY25, ZFCM25]. Construction [YC25]. consumption [CNQY25]. contact [LYC25, XHLH25, ZBN^{+}25]. continuation [LR25]. continuity [Hun25, JJM25]. continuous [DLR25, MRRÁY25, YY25]. continuously [SQLZ25]. contractive [KIF^{+}25]. contractivity [YY25]. contrast [XGYH25]. Control [PSF25, GHZ25, GLTW25, GS25, Kar25b, LS25, MZL25, RNC25, SC25, SKM25, SWWY25, TSM25, XDZ25, ZFCM25]. controllability [Sak25, SX25]. controlled [KIF^{+}25]. controller [JCCZ25]. Convergence [IRS25, LP25b, LMZ25, LR25, Yin25, Hag25, HWZ25, LGDN25, ZHF25]. Convergent [FLS25, SRMGMD25, ZZJJ25]. converging [KS25a]. Convex [AGGM25, HHF25, YPLH25]. convolution [XLZ25]. copositive [HLS25]. core [BPT25, MF25]. correction [MRRÁY25]. corrector [AK25a]. COS [ACO25]. cost [AAN25, FF25]. costs [PW25]. Coupled [LFH25, CG25, JCCZ25, ZMG25]. Cournot [Kar25b]. covering [Hag25]. CP [XNL25]. CP-decomposition [XNL25]. Crank [LCZ25]. credit [HBG25]. cross [XJQ25]. Crouzeix [LZZ25a]. crystal [LLZH25, WX25]. Cubic [LWL^{+}25, GP25]. cumulant [CGR25]. cumulative [KZG25]. curl [WLZZ25]. curves [FKVZ25, PSF25]. CVaR [XLLJ25]. cyclic [BS25c, BRPP25].

D [AKKM25, LMZ25]. damping [ZHF25]. dangling [SHL^{+}25]. Darcy [HJ25, HYZ^{+}25, JLC25, You25].

Darcy-transport [JLC25]. Data [PDT^{+}25, CGR25, DJ25, DLR25, LW25, SQLZ25, WZTL25, WKQ25, WJ25, XDZ25]. Data-driven [PDT^{+}25]. database [HRO^{+}25]. DEA [LG25]. deblurring [FF25]. decentralized [ZLZ25]. decision [CNQY25, Kar25a]. decision-making [CNQY25]. decomposition [CKC25, CWX25, CS25, JS25, KL25, LP25b, LC25b, SGRS25, XNL25]. Decoupled [LWZZ25]. Deep [MR25, SSM25, ZLS25, MR23]. degradation [YSP25]. degree [AM25]. DEIM [CWX25]. delay [GHD25, HRST25, KPTS25, LZZ^{+}25b, PDT^{+}25, XDZ25, Yu25]. delayed [FG25]. delays [JCCZ25, PDT^{+}25, SC25]. delegation [Kar25b]. dendritic [WX25]. dengue [LP25a]. denoising [FF25, ZLC^{+}25]. density [HYZ^{+}25]. dependence [BFQ25]. dependent [CNQY25, OTD^{+}25, WZTL25, YSP25]. derivative [Hag25, LP25a, LJY25]. derivative-free [Hag25, LJY25]. derivatives [GLTW25]. detection [CGR25]. Detective [PM25a]. Determination [QWLF25]. determined [TST25]. determining [KKV25]. Deterministic [GR25, TGDC25]. Developing [HLL25]. DFIG [SC25]. diagnostics [ZLC^{+}25]. diagonal [BS25c, DdF25]. Difference [AGGM25, CTZ25, FG25, FADA25, FC25, HKH^{+}25, JS25, MWTW25, ZZJJ25]. differences [GNUV25]. differencing [HFSK25]. differentiable [SQLZ25]. differential [AFA25, AD25, FADA25, GNUV25, GBB25, HWZ25, HY25b, KSC25, LLG25, PDT^{+}25, Ran25, RM25, SX25, SKM25, XLZ25, Yu25, ZZX25]. diffusion [ACO25, FF25, FGGdO25, FBM25, PY25, XGYH25]. Dilated [XLZ25]. Dimension [XJQ25]. dimensional [Che25, KKHK25, MO25, NKN25, Ran25, ZMG25]. dimensions [ADJL25]. discontinuous [Che25, JLC25, LW25, PM25a]. discovery [PDT^{+}25]. Discrete [GJ25, ADJL25, ACO25, CCLH25, FG25, GH25, Kar25b, LZFP25, LZZ^{+}25b, MHWY25, MWTW25, PDT^{+}25, Sak25, SC25, Vas25]. discrete-time [CCLH25, FG25, MHWY25]. discretization [Buo25, HYZ^{+}25, KL25, SWWY25, SSM25, ZLS25]. discretization-invariant [ZLS25]. Dispersion [SBAM25]. displacement [BFQ25, FLS25]. distance [Xie25]. Distributed [LSCZ25, HRST25]. distributed-order [HRST25]. distribution [CP25, JFP25, KZG25, NN25, YWY25]. distributions [ACO25, TA25, WZTL25]. Dolichobrachistochrone [GBB25]. Domain [KS25a, KL25, LP25b]. domains [YWW25, ZMG25]. Double [CXLP25]. driven [CTZ25, PDT^{+}25, Ran25, ZZX25]. dual [AK25a, CQ25, LR25, ZHF25]. duopoly [Kar25b]. Dynamic [CY25, GBB25, LP25a, XC25]. dynamical [Buo25, ZHF25]. Dynamics [FG25, Yin25, ADY25, MO25].

each [GJ25]. Editorial

[Ano25h, Ano25i, Ano25j, Ano25k, Ano25l, Ano25m, Ano25n, Ano25o, Ano25p]. efficiency [LG25, LZZ25a]. Efficient [CM25, FBZ25, ZNL25, AK25a, CLJX25, CLY25a, GHZ25, GHD25, LC25a, SGRS25, SWWY25, WX25, WKQ25, YWW25, ZLC^{+}25, ZHLD25]. efficiently [SHL^{+}25]. EIEQ [ZSYZ25]. eigenpairs [Yao25]. eigensolutions [ABB25]. eigenvalue [HLS25, Sah25, TC25]. eigenvalues [DdF25, QWLF25, Sah25]. elastic [LWL^{+}25, LGDN25, ZBN^{+}25]. elastodynamic [FLS25]. Electrical [LW25]. electromagnetic [LFH25]. electrophysiology [RNC25]. Element [Vas25, XGYH25, BS25a, CA25b, Che25, CH25, CG25, CKL^{+}25, GHZ25, HWZ25, LCZ25, LYC25, LZZ25a, LMZ25, LSPV25, OTD^{+}25, Sak25, TZ25a, WLZZ25, WCW25, XGX25, ZBN^{+}25]. element-free [LCZ25]. elements [NT25, SWWY25]. ellipse [SH25]. elliptic [ABB25, BS25a, GHZ25, GS25, SWWY25, YWW25, ZH25]. embedding [WJ25]. Emden [SGRS25]. empirical [GH25]. Energy [ZMG25, LYC25, LLZH25, LWZZ25, MCL25, XC25, ZS25, ZSYZ25, ZC25a]. energy- [ZC25a]. Energy-conserving [ZMG25]. energy-optimized [MCL25]. energy-stable [LLZH25]. enhanced [AK25b, FGGdO25]. Enhancing [CA25a]. enriched [GS25]. entropy [LG25]. environment [AAN25]. environments [Hun25]. epidemic [KS25a]. equation [AFA25, AKKM25, BKST25, BCD25, COPPS25, CKL^{+}25, CTZ25, HKH^{+}25, JS25, KKHK25, KIF^{+}25, KM25, LCZ25, LLZH25, LLXB25, LF25, NZ25, RK25, SSM25, XC25, XGX25, Yao25, ZSYZ25, ZZJJ25, ZZ25]. equations [AD25, BPT25, CH25, CG25, FADA25, FBM25, GNUV25, GHD25, HWZ25, HRST25, JJM25, KKV25, KSC25, LFH25, LC25a, LZZ25a, LLG25, LJY25, MS25, MO25, PDT^{+}25, PY25, PM25b, QWZ25, Ran25, RM25, SX25, SQLZ25, SKM25, SGRS25, SBAM25, TZ25a, TZ25b, XLZ25, YTT25, YWW25, Yu25, ZZX25]. equilibria [HY25b]. equilibrium [CXLP25, PY25]. equilibrium-diffusion [PY25]. Erratum [MR25]. Error [GS25, SH25, WCW25, ZSYZ25, AMRS25, CA25a, CA25b, CG25, LTZ25, MWTW25, MR23, MR25, SWD25, TC25, Xie25, YTT25, ZC25a]. estimates [AMRS25, CA25a, CA25b, GS25, MWTW25, MR23, MR25, SWD25, YTT25, ZC25a]. Estimating [BC25]. estimation [FBM25, LLG25]. estimators [NN25]. Euler [Ran25, SKM25]. Eulerian [ADJL25, PY25]. evaluation [LC25b]. even [DLR25]. event [JCCZ25]. event-triggered [JCCZ25]. evolutionary [XHLH25]. exactly [HHF25]. example [OTD^{+}25, ZFCM25]. exchange [Zho25]. exciting [ACO25]. existence [TSM25]. expansion [CA25b]. expansions [FLS25]. explicit [Hag25, KS25a, Ran25, ZSYZ25, ZZX25]. explicit-invariant [ZSYZ25]. exponent [MS25]. exponential [CCLH25, CM25, HFSK25, JCCZ25, KS25c, NPA^{+}25]. exponential-type [KS25c]. extending [CLQY25]. extension [ACO25, CQ25, KS25b, MF25, SRMGMD25, ZLS25]. exterior [QWLF25, Sak25]. extragradient [CXLP25, OAO25]. extrapolation [LC25b].

facility [LZFP25, MHWY25]. factor [LLG25]. factorization [ZNL25]. factors [LP25a]. Fading [BCD25]. failure [DJ25, GJ25]. failures [SDS25, YSP25]. family [WZTL25]. Fast [ZHF25, LCZ25]. FDTD [HLL25]. FE [ZC25a]. February [Ano25q]. feedforward [YC25]. Feller [CdlHG25]. Fickian [FGGdO25]. field [LLZH25, LWZZ25, MT25, WX25, Yin25]. Fields [BAPPF25]. filled [SQLZ25]. finance [HBG25]. find [SSLP25]. finding [FA25]. Finite [CTZ25, FC25, KPTS25, SZD25, Vas25, XGYH25, BS25a, CA25b, CLZ25, Che25, CH25, CKL^{+}25, FG25, FADA25, GNUV25, GHZ25, HKH^{+}25, HWZ25, LYC25, LMZ25, LSPV25, MWTW25, NT25, PY25, Sak25, SWWY25, TZ25a, XGX25, ZBN^{+}25]. Finite-frequency [SZD25]. finite-part [CLZ25]. first [Buo25, CA25b, XC25]. first-order [CA25b, XC25]. Fisher [KKHK25]. fitting [SQLZ25]. fixed [ADY25, CXLP25, LSCZ25, MZL25]. fixed-time [MZL25]. flow [Che25, CLY25b, GCN25, HYZ^{+}25, UX25, You25]. fluid [ADY25]. Fokker [RK25]. footrule [TST25]. Forchheimer [HYZ^{+}25]. formation [FBM25]. formula [CA25a, CA25b]. formulas [SHK25]. formulation [ZZ25]. Forward [YXM25, CP25, IRS25]. four [CM25]. fourth [HKH^{+}25, LMZ25, SGRS25]. fourth-order [HKH^{+}25, LMZ25, SGRS25]. Fowler [SGRS25]. FPMPE [ZWSC25]. fractal [KCM25]. fractional [ADY25, ABH^{+}25, AAN25, CH25, CG25, GNUV25, GLTW25, HFSK25, HRST25, HuR25, JN25, KPTS25, Kar25b, KSC25, LP25a, Öza25, RK25, RM25, SRMGMD25, SKM25, TZ25a, XDZ25, ZLC^{+}25, ZZX25]. fractional-order [GLTW25, XDZ25, ZLC^{+}25]. fracture [Che25]. fractured [ABH^{+}25]. fragile [XDZ25]. framework [GR25, Kar25a, KS25b]. Fréchet [NN25]. free [Hag25, LCZ25, LJY25, SQLZ25, WJ25]. frequency [SZD25]. fully [MWTW25]. function [CGR25, KZG25, MHO25, MPB25, Öza25, Pho25, SQLZ25, ZC25b]. functional [COPPS25]. functionals [KKV25]. functions [CLZ25, FLS25, KS25c, KCM25, LSPV25, LC25b, MRRÁY25, OTD^{+}25, SHK25, SH25, YC25]. fundamental [KC25]. Fuzzy [KS25c, AAN25, Hun25, SC25].

Galerkin [Che25, JLC25, LCZ25, LW25, LMZ25, PM25a, PM25b, QWZ25, ZMG25]. game [GBB25, Kar25b]. games [HY25b, Hun25]. gap [QWLF25]. Gauss [SH25]. Gaussian [BAPPF25, Ran25]. general [HuR25, MRRÁY25, WCW25, Yu25]. Generalized [HFSK25, Vas25, HY25b, Hun25, KM25, LC25a, LYC25, MG25, MO25, NPA^{+}25, PM25a, Ran25, SWD25, WZTL25, XGX25, YWY25, ZZJJ25, ZZ25, XGYH25]. generating [CGR25]. generative [BDvdHV25]. Generic [MHO25]. genuine [CQ25]. Geometric [AAN25, CLY25a, DJ25]. Gibson [BFQ25]. Ginzburg [LCZ25]. Global [SWD25, FA25, LSCZ25]. globally [YY25]. GMsFEM [SSM25]. GMsHDG [PM25a]. Goal [BS25b]. Goal-oriented [BS25b]. golden [LY25, OAO25]. Gompertz [JFP25]. Gordon [ZMG25]. governed [FBM25]. grade [UX25]. gradient [AMRS25, CLQY25, GLTW25, JWS^{+}25, ZHLD25]. gradient-based [GLTW25]. Gramians [XJQ25]. graph [SNZ25, Vas25]. graphs [BAPPF25, LSCZ25]. gravitational [Sai25]. greedy [ABB25, LLXB25]. Gregory [ZBK25]. grid [CH25, GCN25, LRG25, TZ25a, WCW25]. growth [WX25].

Hamiltonian [BS25b]. Hankel [FBZ25]. harmonic [CS25, Sak25]. Hawkes [ACO25]. health [CNQY25]. heat [ADY25, HYZ^{+}25, NKN25, OTD^{+}25]. Helmholtz [CTZ25]. Hemispheroidal [CS25]. hemivariational [XHLH25]. Hermite [ZMG25]. Hermitian [Pho25, XNL25]. Hessenberg [BKST25]. Heston [ACO25, HKY25]. heterogeneous [ABH^{+}25, DJ25, GCN25, UX25]. heuristic [ZS25]. HH [BRPP25]. High [RK25, SBAM25, BAPPF25, CWX25, WX25, XGYH25]. high-contrast [XGYH25]. High-order [RK25, SBAM25, WX25]. Higher [LGDN25, XNL25]. Higher-order [LGDN25]. Highly [SGRS25, AKKM25, CLZ25, ZC25b]. Hilbert [SHK25]. history [BFQ25]. hodograph [FKVZ25, PSF25]. homogeneous [UX25]. HOMS [LGDN25]. HOMS-PINN [LGDN25]. horizontal [MG25, WCS25]. hybrid [Che25, JWS^{+}25, YPLH25, Zho25]. hybrid-dimensional [Che25]. hybridizable [PM25a]. hydrodynamics [PY25]. hyperbolic [YWY25]. hypergeometric [Pho25]. hyperplane [WYZ25]. identification [LZZ^{+}25b]. identity [BRPP25]. IDFPM [JJM25]. IDFPM-based [JJM25]. image [CKC25, CXLP25, HRO^{+}25, IRS25, JJM25, ZLC^{+}25]. image-based [ZLC^{+}25]. images [HRO^{+}25]. imaging [MT25]. IMEX [YTT25]. immersed [ZH25]. impact [DG25]. Impedance [LW25]. imperfect [SDS25]. implementation [BPT25]. implicit [LS25, WCS25]. imprecise [XQW25]. improve [HRO^{+}25]. Improved [WCS25, HBG25, OAO25]. Improving [MCL25]. impulses [CCLH25]. inclusion [HLS25, IRS25]. independent [LZZ25a]. index [JFP25]. indicator [Kar25a]. indicator-based [Kar25a]. induced [CWX25]. induction [PM25b]. inequalities [Buo25, BRPP25, LY25, TSM25, XHLH25]. inequality [OAO25]. Inertial [HHF25, CXLP25, IRS25, JCCZ25, LJY25, ZHF25]. inexact [CLY25b, ZLZ25]. inexactly [HHF25]. infeasible [AK25a]. inference [DJ25, KZG25, SDS25]. infinite [DLR25, LC25b]. Influence [Sai25]. information [BRPP25]. informed [LGDN25]. inhomogeneous [TC25]. initial [Ran25]. inner [MHO25]. Insights [CNQY25]. instabilities [FA25]. instability [SBAM25]. insurance [TA25]. integer [HY25a]. integer-valued [HY25a]. integrable [MO25]. Integral [MPB25, FLS25, GHZ25, GHD25, KIF^{+}25, LFH25, MS25, ZZ25]. integrals [CLZ25, LC25b, SHK25, ZC25b]. integrated [Kar25a]. integration [KS25b, MS25, MRRÁY25, NPA^{+}25]. integrators [LWF25, XGYH25]. integro [AFA25, AD25, FADA25, RM25, SKM25, Yu25]. integro-differential [AFA25, AD25, FADA25, RM25, SKM25, Yu25]. interest [Zho25]. interface [BS25a, LP25b, ZH25]. interior [GS25]. interpolation [CA25a, CA25b, GH25, MHO25, OTD^{+}25]. interpolatory [CWX25]. Intrusive [GCN25]. Invariant [QWZ25, ZLS25, ZSYZ25]. invariants [LWF25]. inventory [CY25]. Inverse [Sah25, BCD25, CQ25, FBM25, MHO25, MF25, NKN25, YXM25]. Inversion [YXM25]. inverted [WZTL25]. investigate [FA25]. investment [PW25]. involving [LC25b]. IPM [AK25a]. irregular [YWW25]. Ishikawa [LSCZ25]. isogeometric [LSPV25]. isotropic [TC25]. iteration [AK25a, WCS25, ZWSC25]. iterative [GHZ25, LYC25, MHO25, ZLY25].

J [MR25]. January [Ano25a, Ano25d]. Jensen [BRPP25]. jobs [SSLP25, ZS25]. Joint [FBM25, YWY25]. judgment [KZG25]. jump [ACO25]. jump-diffusion [ACO25]. June [Ano25r].

Kac [DdF25]. Kaczmarz

[LLXB25, NZ25, TGDC25, WYZ25]. Kawahara [SRMGMD25]. KdV [ZZJJ25, BPT25]. Keller [FBM25]. kernel [JN25, KS25c, Öza25]. kernels [HuR25, SBAM25]. Kerr [CKL^{+}25]. Kerr-type [CKL^{+}25]. kind [AD25, GHD25, LF25]. kinetic [FA25]. Klein [ZMG25]. known [AK25a, FLS25]. Kolmogorov [KKHK25]. Kudryashov [MO25]. Kutta [Moi25, KS25b].

L [CWX25]. L-DEIM [CWX25]. Lagrange [WKQ25]. Lagrangian [ADJL25, HHF25, PY25, ZNL25]. Lanczos [LTZ25]. Landau [LCZ25]. Lane [SGRS25]. Laplace [MR25, MR23]. large [AK25a, DLR25, TGDC25]. large-scale [TGDC25]. laws [ADJL25, FC25]. layer [ZBN^{+}25]. layers [ADY25]. LCP [SWD25]. learn [NT25]. learning [SSM25]. Least [Ton25]. Leffler [JN25, Öza25]. length [BC25]. level [KC25]. Levenberg [TZ25b]. LFC [SC25]. liabilities [ZFCM25]. life [TA25]. lifetime [JFP25, WZTL25]. lifetimes [DJ25]. LightGBM [HBG25]. Lightning [SNZ25]. like [CA25a, TZ25b]. likelihood [DJ25, LLG25]. limit [PY25]. limited [XJQ25]. linear [CLJX25, GHZ25, HV25, LLZH25, LZZ^{+}25b, MG25, SX25, SZD25, TGDC25, Yu25, ZFCM25]. linearized [CG25]. linearly [HHF25]. Lipschitz [GR25, JJM25, LY25, YY25, ZZX25]. liquidity [LG25]. Lobatto [SH25]. local [FC25, KL25, Xie25, You25]. location [LZFP25]. locations [LZFP25]. Logarithmic [NN25]. logistic [LJY25]. long [WKQ25]. long-only [WKQ25]. loop [HY25b]. lot [SSLP25]. Low [FF25, BKST25, SZD25]. Low-cost [FF25]. low-rank [BKST25, SZD25]. Lucas [LC25b]. Luchko [HuR25].

machine [SSLP25]. magnetic [Yin25]. Mahony [KM25]. making [CNQY25]. management [ZFCM25]. manifolds [JWS^{+}25]. mapping [YWW25]. mappings [KIF^{+}25]. maps [Hun25, SC25]. March [Ano25b, Ano25e]. Markov [BAPPF25]. Marquardt [TZ25b]. mass [ZC25a]. mass- [ZC25a]. matched [OTD^{+}25]. matching [SNZ25]. material [UX25]. materials [LGDN25]. Math [MR25]. mathematical [ITM25, SGRS25]. matrices [BS25c, CQ25, Sah25]. matrix [CKC25, DdF25, FWL25, LC25a, MG25, NZ25, Pho25, WCS25, ZNL25]. maximization [CGR25]. Maximum [DJ25, LLG25, WZTL25]. Maxwell [CKL^{+}25]. May [Ano25f, Ano25c]. Mean [Hag25, PW25]. Mean-square [Hag25]. mean-variance [PW25]. means [XQW25]. measles [ITM25]. measure [SX25]. measures [LG25]. mechanics [XHLH25]. mechanism [MHWY25]. media [ABH^{+}25, CKL^{+}25, GCN25, PM25a, XGYH25, You25]. medical [ZLC^{+}25]. medium [Sai25, TC25]. memristive [JCCZ25]. mesh [LMZ25, You25]. meshes [WCW25]. Meshfree [NPA^{+}25]. meshless [AD25, GNUV25]. metasurfaces [HLL25]. Method [Vas25, ADJL25, ABB25, AMRS25, AFA25, AD25, ACO25, BKST25, BS25a, BCD25, COPPS25, Che25, CG25, FADA25, FWL25, GCN25, GHD25, GS25, Hag25, HKH^{+}25, HJ25, HWZ25, HLL25, JS25, JLC25, KC25, KSC25, LCZ25, LW25, LP25b, LWL^{+}25, LYC25, LRG25, LGDN25, LMZ25, LLXB25, LSPV25, LF25, MS25, MCL25, MR23, MR25, NKN25, NN25, NPA^{+}25, NDSZ25, OTD^{+}25, PM25a, QWZ25, QWLF25, SRMGMD25, SKM25, TGDC25, TZ25a, WLZZ25, WYZ25, WCW25, XNL25, XGX25, XQW25, YTT25, YWW25, Yao25, ZLY25, ZMG25, ZBN^{+}25, ZWSC25, ZNL25]. Methods [XGYH25, AS25, CM25, CH25, CKL^{+}25, CTZ25, GH25, HV25, KS25b, LC25b, MWTW25, Moi25, PM25b, WX25, WCS25, Yin25, Yu25, ZZX25]. metric [FKVZ25, KIF^{+}25]. MFPIO [ZWSC25]. Milstein [YY25]. Milstein-type [YY25]. minimization [LSPV25]. minimizers [Xie25]. minimizing [LYC25]. Mittag [JN25, Öza25]. Mittag-Leffler [JN25, Öza25]. Mixed [ZBN^{+}25, Che25, LY25]. mKDV [BPT25]. mobile [MZL25]. mobility [LWZZ25]. Model [WJ25, ABH^{+}25, ACO25, CY25, CLY25a, FG25, HKY25, ITM25, Kar25b, KS25a, LP25a, LRG25, LWZZ25, RNC25, SZD25, WX25, XLLJ25, ZLZ25, ZZ25]. modeling [MHWY25, YXM25]. modelling [BDvdHV25]. models [Che25, CNQY25, HFSK25, LG25, Vas25, Zho25]. modification [FKVZ25]. modifications [PSF25]. modified [BKST25, CKL^{+}25, KKHK25, KM25, QWLF25, ZHLD25]. Modulus [MG25, FWL25, WCS25]. Modulus-based [MG25, FWL25, WCS25]. moment [Ton25]. moments [CP25, NN25, YWY25]. momentum [LLXB25, TGDC25]. monodomain [RNC25]. monotone [CLQY25, IRS25, LJY25, XLLJ25]. Monte [SDS25]. Montgomery [BRPP25]. Moore [BFQ25, CQ25]. MOR [ABB25]. Moreau [RNC25]. motion [AAN25, ZZX25]. MP [Yao25]. MP-Newton [Yao25]. Multi [KC25, BPT25, KSC25, LGDN25, LSCZ25, LLG25, SHL^{+}25, ZMG25, ZBN^{+}25]. multi-agent [LSCZ25]. multi-core [BPT25]. multi-dimensional [ZMG25]. multi-factor [LLG25]. multi-layer [ZBN^{+}25]. Multi-level [KC25]. multi-scale [LGDN25]. multi-step [SHL^{+}25]. multi-term [KSC25]. multilevel [XGX25, You25]. multilinear [NDSZ25, ZWSC25]. multiobjective [CLQY25, Hun25]. Multiscale [ABH^{+}25, Vas25, XGYH25, LYC25, NPA^{+}25, PM25a, XLZ25]. multiterm [KPTS25]. multivariate [CGR25, Öza25, Xie25, YWY25].

n [GJ25]. Nash [HY25b]. nature [PSF25]. near [MT25]. near-field [MT25]. network [LGDN25, MZL25, Vas25, YC25, ZH25]. networks [JCCZ25, NT25, ZLS25]. Neural [YXM25, JCCZ25, LGDN25, MZL25, NT25, XLZ25, YC25, ZH25]. neutral [AFA25, FADA25, KPTS25, Yu25]. Newton [CLY25b, HV25, LC25a, WKQ25, Yao25]. Newton-type [HV25, LC25a]. Nicolson [LCZ25]. noise [CTZ25, Hag25, YPLH25]. noises [Ran25]. Non [DG25, FGGdO25, GCN25, XDZ25, CLQY25, FC25, Hag25, LY25, RM25, SX25, YY25, YPLH25]. non-commutative [Hag25]. non-convex [YPLH25]. Non-Fickian [FGGdO25]. Non-fragile [XDZ25]. non-globally [YY25]. Non-Intrusive [GCN25]. non-linear [SX25]. non-Lipschitz [LY25]. non-monotone [CLQY25]. non-smooth [RM25]. Non-uniformly [DG25]. nonconvex [CLJX25, JWS^{+}25, Xie25]. nonlinear [CH25, CG25, CKL^{+}25, FF25, FWL25, Hag25, HRST25, JJM25, KL25, LCZ25, LRG25, LLXB25, LJY25, MO25, SQLZ25, TZ25b, WCW25, XGX25, YTT25, Yao25, Yu25, ZMG25, ZC25a]. nonlocal [ADJL25, LP25b, Sai25]. nonnegative [ZNL25]. nonsmooth [CLY25b, CLQY25]. nonstandard [FG25]. nonzero [HY25b]. nonzero-sum [HY25b]. norm [LSPV25, Xie25]. note [DdF25]. Novel [GHZ25, CA25a, FF25, HWZ25, JCCZ25, MZL25, XGX25]. nudging [DLR25]. numbers [GJ25]. Numerical [AD25, BFQ25, BPT25, CdlHG25, FADA25, HYZ^{+}25, KM25, YY25, ZBK25, AFA25, GNUV25, JVW25, JN25, KKHK25, MRRÁY25, RK25, ZSYZ25, ZFCM25]. Nuttall [MPB25]. Nyström [MS25].

objective [LG25]. observations

[BC25, XQW25]. ODE [LS25]. off [DdF25]. off-diagonal [DdF25]. one [NKN25, ZZX25]. one-dimensional [NKN25]. one-sided [ZZX25]. online [SSM25]. only [WKQ25]. open [CS25, HY25b]. open-loop [HY25b]. Operator [YXM25, LWZZ25, XLZ25, ZLS25]. operators [CdlHG25, JN25, LSCZ25, Öza25, YC25]. Optical [MO25]. Optimal [MT25, SHK25, ZC25a, GHZ25, GLTW25, GS25, LS25, LSPV25, RNC25, SSLP25, SKM25, SGRS25, SWWY25, ZLC^{+}25, ZFCM25]. Optimally [AKKM25]. optimization [CLY25a, CLQY25, GLTW25, HHF25, JWS^{+}25, LR25, MHWY25, ZLZ25]. optimization-by-continuation [LR25]. optimized [HBG25, MCL25]. option [AAN25, MWTW25]. options [ACO25, HKY25, ZZ25, Zho25]. order [AK25a, AFA25, Buo25, CM25, CWX25, CA25b, DG25, FADA25, GLTW25, HKH^{+}25, HRST25, KPTS25, Kar25b, LLZH25, LRG25, LGDN25, LMZ25, LWF25, LC25b, Moi25, Ran25, RK25, SGRS25, SZD25, SBAM25, TZ25a, WX25, XC25, XNL25, XDZ25, ZLC^{+}25, ZZJJ25]. order-book [DG25]. oriented [BS25b]. Orr [PM25b]. orthogonal [JS25]. oscillating [SHK25]. oscillation [HFSK25]. oscillatory [AKKM25, CLZ25, LWF25, ZC25b]. Outlier [CGR25].

padded [BS25c]. PageRank [SHL^{+}25, ZWSC25]. parabolic [NPA^{+}25, QWZ25, SBAM25, WCW25, YTT25]. Parallel [BS25c, SSLP25, ZS25]. parallel-batch [SSLP25]. parallelization [BPT25]. parameter [FBM25, SQLZ25]. parameter-free [SQLZ25]. parameterization [CS25]. parameterized [AK25b]. parameters [DJ25, DLR25]. parametric [ABB25, Hun25]. Pareto [HLS25, SSLP25]. part [CLZ25]. partial [HWZ25, LW25, XLZ25]. Particle [Yin25]. path [AK25a, MZL25]. paths [BAPPF25]. pattern [FBM25, KKV25]. PDE [ZLC^{+}25]. PDEs [ABB25, FBM25]. Peano [JN25]. Peer [LS25]. penalized [XLLJ25]. penalty [GS25, HJ25]. Penrose [CQ25]. performance [JFP25]. periodic [LZZ^{+}25b, SHK25]. perishable [CY25]. person [HY25b]. perturbations [DdF25]. perturbed [FBZ25, LMZ25]. Petrov [PM25b]. Petrovsky [KKHK25]. phase [Che25, LLZH25, LWZZ25, WX25, ZFCM25]. phase-field [LWZZ25, WX25]. physics [LGDN25, TSM25]. physics-informed [LGDN25]. piece [MRRÁY25]. piece-wise [MRRÁY25]. PINN [LGDN25]. Piskunov [KKHK25]. planar [FKVZ25, PSF25]. Planck [RK25]. plasma [FA25]. plus [FBZ25]. POD [LRG25]. point [ADY25, AK25b, CXLP25, PSF25]. points [LSCZ25]. pointwise [RNC25]. Poisson [YPLH25]. poles [FA25]. polygonal [WCW25]. polyharmonic [KC25]. polyhedral [SWD25]. Polynomial [SSLP25, FC25, Sah25]. polynomials [FKVZ25, NKN25, PM25b]. Population [LZFP25]. Population-based [LZFP25]. poroelasticity [LRG25]. porous [GCN25, PM25a, You25]. port [BS25b]. port-Hamiltonian [BS25b]. portfolio [CNQY25, LG25, WKQ25]. posedness [XHLH25]. Positivity [KKHK25, PY25]. positivity-preserving [PY25]. post [HWZ25, KZG25]. post-processed [HWZ25]. posteriori [MWTW25, MR23, MR25, YTT25, You25]. posteriori-based [You25]. Powell [GP25]. power [AAN25, CLY25b, XNL25]. PR [CLJX25]. preconditioner [AK25b]. preconditioning [MHO25]. predator [FG25]. Prediction [SSM25, HBG25]. predictor [AK25a]. predictor-corrector [AK25a]. prescribed [Sah25]. preservation [BDvdHV25]. preserve [LWF25, PSF25]. preserving [KKHK25, KS25a, PY25]. prey [FG25]. price [CY25, DG25]. Pricing [HKY25, Zho25, AAN25, ACO25, CY25, MWTW25, ZZ25]. primal [AK25a, LR25, ZHF25]. primal-dual [AK25a, LR25, ZHF25]. priori [MR23, MR25]. problem [AS25, ABH^{+}25, BFQ25, BCD25, CLY25b, FWL25, GS25, HJ25, JLC25, LYC25, LMZ25, LF25, MR23, MR25, OTD^{+}25, Sah25, TC25, WYZ25, XLLJ25, You25]. problems [AK25a, AK25b, BS25a, CLJX25, CLY25a, CLY25b, CLQY25, CXLP25, FBM25, GCN25, GHZ25, GLTW25, HHF25, HV25, IRS25, JJM25, JCCZ25, JWS^{+}25, KC25, LC25a, LGDN25, LJY25, LY25, MG25, NKN25, NPA^{+}25, OAO25, SQLZ25, SSLP25, SKM25, SWWY25, WLZZ25, WCS25, WCW25, XNL25, YXM25, ZH25]. procedure [WZTL25]. process [ACO25, CP25, HY25a]. processed [HWZ25]. product [LC25b, MS25, MHO25]. products [CY25]. profit [Kar25b]. programming [AGGM25, CLY25a, CLY25b, GBB25, LG25, XLLJ25]. programs [CLJX25]. projection [LJY25]. Projective [KS25b]. propagation [LWL^{+}25, Sak25]. proper [JS25]. property [QWZ25]. proportional [Kar25a]. proportional-XL [Kar25a]. proposed [FG25]. proximal [CLQY25]. pseudo [LJY25]. pseudo-monotone [LJY25]. pulmonary [HRO^{+}25]. pursuit [WYZ25]. Pythagorean [FKVZ25, PSF25]. Pythagorean-hodograph [FKVZ25, PSF25].

quad [WLZZ25]. quad-curl [WLZZ25]. quadratic [JVW25, Xie25]. quadratization [MCL25, ZSYZ25]. quadrature [SHK25, SH25, ZBK25, ZC25b]. quality [CY25]. quantiles [GR25]. quasi [JCCZ25]. quasi-synchronization [JCCZ25]. Queue [ACO25, BC25]. quintic [PSF25]. quotients [JS25].

radial [KCM25, MHO25]. radiation [PY25]. Random [BAPPF25, SWWY25]. Randomized [LLXB25, NZ25]. range [JVW25]. rank [BKST25, SZD25, XNL25]. rank- [XNL25]. ranked [WZTL25]. ranking [LZFP25]. rate [YC25, Zho25]. rates [BC25, ZHF25]. ratio [LY25, OAO25]. Rational [GP25]. rationality [HY25b]. Raviart [LZZ25a]. Rayleigh [TZ25a]. reaction [FBM25]. reaction-advection-diffusion [FBM25]. reactions [UX25]. real [WKQ25]. Reciprocal [TSM25]. reciprocity [QWLF25]. record [JFP25]. recovery [IRS25]. rectangular [KIF^{+}25]. recurrence [CP25]. recursive [LZZ^{+}25b]. Reduced [GCN25, CKC25, GS25, LRG25]. reduced-order [LRG25]. reduction [ABH^{+}25, BS25c, SZD25, XJQ25, Xie25]. reference [CY25]. refined [CA25b]. refinement [You25]. reflected [IRS25]. reflected-forward-backward [IRS25]. regime [AKKM25, ZZ25]. regime-switching [ZZ25]. region [QWZ25, TST25, Xie25]. regression [KS25c, LC25a, LJY25, Ton25]. regularization [BCD25, LF25, RNC25]. regularized [CLJX25, WYZ25, ZHF25]. regularizers [YPLH25]. reinsurance [Kar25a, PW25]. relative [BAPPF25, Kar25b]. Relaxation [LWF25]. relaxed [LJY25]. relaxed-inertial [LJY25]. Reliability [SDS25, YSP25, LZZ25a]. removal [YPLH25]. renewal [CP25]. reordering [SHL^{+}25]. reorthogonalization [LTZ25]. repair [MHWY25, SDS25]. repairable [MHWY25]. Representation [Pho25, AS25, GP25]. representations [MPB25]. residual [LSPV25]. resolution [HRO^{+}25]. resource [LY25]. respect [MHO25]. restart [JWS^{+}25]. restoration [CKC25, CXLP25, JJM25]. results [RM25, TSM25]. retrial [MHWY25]. Revisiting [TST25]. Richards [SSM25]. Richardson [NDSZ25]. ridge [LC25a]. Riemannian [JWS^{+}25]. Riesz [CdlHG25]. risk [AAN25, CLY25a, HBG25]. risks [WZTL25]. risky [LG25]. Ritz [MR25, MR23]. RKN [LWF25]. RKN-type [LWF25]. RLW [ZZJJ25]. robots [MZL25]. Robust [SC25, AK25b, CLQY25, KS25c, PW25]. role [CNQY25, ITM25]. Rosenau [SRMGMD25, ZZJJ25]. Rosenbrock [CM25]. rotating [Sai25]. ruled [GP25]. rules [MRRÁY25]. Runge [KS25b, Moi25].

S [SC25]. SAA [CLY25b]. SAA-based [CLY25b]. Sabin [GP25]. saddle [AK25b]. sample [XLLJ25]. sampled [DG25, XDZ25]. sampled-data [XDZ25]. samples [WZTL25]. sampling [WZTL25]. Sard [JN25]. SAT [KM25]. SAV [ZMG25]. SBP [KM25]. SBP-SAT [KM25]. scale [LGDN25, TGDC25]. scaling [HHF25, ZS25, ZHF25]. scattering [LFH25]. schedules [SSLP25]. scheduling [SSLP25, ZS25]. scheme [FG25, HRST25, KKHK25, KS25a, LLZH25, PY25, Ran25, XC25, ZSYZ25, ZMG25, ZNL25]. schemes [FC25, GHZ25, LWZZ25, RK25, YY25, ZC25a, ZZJJ25]. Schrödinger [AKKM25, CH25, CG25, LF25, MO25, XGX25, Yao25, ZC25a]. Schrödinger-type [ZC25a]. Schur [TA25]. Schur-constant [TA25]. sciences [COPPS25]. SDEs [Hag25, YY25]. SDO [AK25a]. search [HBG25]. seasonal [LP25a]. Second [AFA25, AK25a, FADA25, JS25, LLZH25, LWF25, Ran25, TZ25a]. Second-order [AFA25, AK25a, FADA25, LLZH25, LWF25, Ran25, TZ25a]. sections [OTD^{+}25]. seeking [LSCZ25]. Segel [FBM25]. selection [WKQ25]. Self [AS25, ACO25, OAO25]. self-adaptive [OAO25]. self-exciting [ACO25]. Semi [ADJL25, Hun25, MHO25]. semi-continuity [Hun25]. Semi-discrete [ADJL25]. semi-inner [MHO25]. semilinear [QWZ25, TZ25a, Yao25]. semismooth [CLY25b]. sensing [CLJX25, ZHLD25]. sensor [MT25]. sequence [BPT25]. series [MPB25]. service [BC25]. set [WZTL25]. sets [WKQ25]. several [ADJL25]. share [LY25]. shift [AK25b]. shift-splitting [AK25b]. shifted [XNL25]. Shishkin [LMZ25]. sided [ZZX25]. sigmoidal [YC25]. signal [JJM25]. simplex [ZC25b]. simply [CS25]. simulated [DG25, HRO^{+}25]. simulation [HLL25, KM25, SDS25]. single [SSLP25]. singular [CKC25, MS25]. singularly [LMZ25]. SIS [KS25a]. sixth [ZZJJ25]. sixth-order [ZZJJ25]. Smart [AM25]. SMEs [HBG25]. smooth [RM25]. SOC [CLY25b]. solitons [MO25]. Solution [KIF^{+}25, GNUV25, GBB25, Hun25, KSC25, TSM25]. solutions [BPT25, FADA25, KC25, RM25, SHL^{+}25]. solve [GLTW25, HHF25, SRMGMD25]. solvers [MHO25]. Solving [Yu25, AK25a, AD25, CLQY25, CXLP25, GHD25, HV25, HRST25, IRS25, KC25, LC25a, LGDN25, LJY25, LF25, NZ25, NDSZ25, RK25, SQLZ25, SGRS25, TGDC25, TZ25b, XNL25, Yao25, ZSYZ25]. Some [TA25, Yin25, ZLS25]. Sommerfeld [PM25b]. sources [AAN25]. space [BS25a, CH25, GNUV25, SHK25]. space-fractional [CH25]. spaces [KIF^{+}25, OAO25]. sparrow [HBG25]. sparse [JJM25, Sah25, WYZ25]. Spearman [TST25]. Spectral [TC25, GHD25, HRST25, PM25b, WLZZ25, ZMG25]. speed [ZS25]. SPH [SBAM25]. spline [LWL^{+}25, RK25]. splines [GP25]. split [YY25]. split-step [YY25]. splitable [SSLP25]. splitting [AK25b, CLJX25, FWL25, IRS25, LC25a, LWZZ25, MG25, WCS25, Yin25]. square [Hag25]. squares [Ton25]. Stability [FADA25, HY25b, Kar25b, CCLH25, KPTS25, SX25]. Stabilization [LSPV25, CCLH25]. stabilized [LZZ25a]. stable [CKL^{+}25, KKHK25, LLZH25, LWZZ25, XC25, ZZJJ25]. standard [LJY25]. standby [MHWY25]. state [CNQY25, FBM25, GHZ25, GJ25, GS25, RNC25]. state-dependent [CNQY25]. state-parameter [FBM25]. stationary [AKKM25, HJ25]. Statistical [KZG25]. Stefan [NKN25]. Steklov [QWLF25]. step [LJY25, Moi25, SRMGMD25, SHL^{+}25, YY25]. stepping [WX25]. stepsize [LS25]. Stochastic [MHWY25, CCLH25, CLY25b, CTZ25, HKY25, KS25a, LY25, XLLJ25, ZLZ25, ZZX25, Zho25]. Stokes [MR25, HJ25, KL25, MR23, TZ25a]. Strang [Yin25]. strategies [PW25]. strategy [JWS^{+}25, ZLC^{+}25]. stratification [KZG25]. strong [Yin25]. strongly [KS25a, SHK25]. structural [ZLY25]. structure [CLY25a]. structured [Sah25]. study [PM25b, UX25]. subdiffusion [YXM25]. subgradient [CXLP25, OAO25]. Sufficient [Xie25]. sum [HY25b]. Superconvergent [RM25]. supply [HBG25]. sure [CCLH25]. surfaces [AM25, CS25, GP25]. surfactant [LWZZ25]. surrogate [WYZ25]. SVM [AGGM25]. switching [SDS25, ZZ25]. Sylvester [BKST25, DdF25]. symmetric [ZNL25]. symmetrical [FBZ25]. synchronization [JCCZ25]. system [Buo25, DJ25, GJ25, KL25, SRMGMD25, YSP25, ZMG25, ZC25a, ZHF25]. systems [BS25b, CCLH25, FBZ25, KPTS25, LWF25, LZZ^{+}25b, MHWY25, NDSZ25, SZD25, TGDC25, XJQ25, XDZ25, ZLY25, ZBN^{+}25].

T [SC25]. Tail [YWY25]. Takens [WJ25]. Taylor [CA25a]. Taylor-like [CA25a]. technique [AD25, FF25, LWZZ25, MCL25, MO25, OAO25, SWWY25, TZ25b]. techniques [HRO^{+}25]. temperature [FGGdO25]. tempered [HuR25]. Temporal [TZ25a]. tensile [SBAM25]. tensor [AS25, BKST25, CWX25, NDSZ25, XLLJ25]. tensors [HLS25, NDSZ25]. term [KSC25]. terms [FLS25]. test [LSPV25]. testing [Kar25b]. Thailand [LP25a]. theorem [JN25]. theorems [HLS25]. theory [LG25]. thermal [BFQ25, Sai25]. thermoelastic [BFQ25, Sai25]. thermoporoelasticity [ABH^{+}25]. Third [Moi25, AD25, GHD25, UX25]. third-grade [UX25]. Thompson [BFQ25]. three [AK25b, GJ25, KKHK25, LC25b]. three-by-three [AK25b]. three-dimensional [KKHK25]. three-state [GJ25]. Tikhonov [ZHF25]. till [CM25, DJ25]. time [ABH^{+}25, BS25b, CCLH25, CG25, FG25, GJ25, HFSK25, JCCZ25, KPTS25, KM25, KS25b, KL25, LSCZ25, LZZ^{+}25b, MHWY25, MZL25, RK25, Sak25, SSLP25, SC25, TZ25a, WX25, XJQ25, XGYH25, XDZ25, ZMG25, ZHF25]. time-delay [LZZ^{+}25b]. time-fractional [CG25, RK25, TZ25a]. time-harmonic [Sak25]. time-limited [XJQ25]. time-stepping [WX25]. time-varying [KPTS25, LSCZ25, SC25, XDZ25]. timer [HKY25]. times [CP25]. Timestep [OTD^{+}25]. Timestep-dependent [OTD^{+}25]. Toeplitz [FBZ25]. Toeplitz-plus-Hankel [FBZ25]. Tomography [LW25]. tracking [ABB25, MZL25]. trajectory [ZHF25]. transaction [AAN25, PW25]. transfer [ADY25, HYZ^{+}25]. transformed [BS25a, CG25]. transmission [ITM25, LP25a, TC25]. transport [GCN25, JLC25]. treatment [ITM25]. triangles [ZBK25]. triangular [WLZZ25]. triggered [JCCZ25]. triplets [LS25]. truncated [AKKM25, Hag25]. truss [CLY25a]. trust [Xie25]. Turkey [ITM25]. Two [CH25, JWS^{+}25, LRG25, ZZJJ25, ZZX25, CLJX25, Che25, GCN25, LWF25, LJY25, Moi25, SRMGMD25, TZ25a, WCW25]. two-block [CLJX25]. Two-grid [CH25, LRG25, GCN25, TZ25a, WCW25]. two-phase [Che25]. two-step [LJY25, Moi25, SRMGMD25]. type [BFQ25, BRPP25, CLJX25, CKL^{+}25, DdF25, FBM25, HV25, KS25c, LC25a, LWF25, SWWY25, SWD25, TZ25b, XC25, YY25, ZC25a].

ultrasound [ZLC^{+}25]. unbounded

[ZMG25]. Uncertain

[LG25, XQW25, CLY25a, LLG25, XDZ25]. uncertainty [LG25]. Unconditional [CG25, CKL^{+}25]. unconditionally [KKHK25, LLZH25]. undamped [ZLY25]. unequal [WZTL25]. uniform [MT25]. uniformly [DG25]. unreliable [MHWY25]. update [TZ25b]. updates [AK25a]. updating [SHL^{+}25, ZLY25]. Using [BPT25, AD25, CA25a, FF25, GNUV25, HRO^{+}25, KM25, KZG25, MO25, SSM25, XLLJ25]. utilizing [PM25b]. vaccination [ITM25]. value [BCD25, CKC25, CLY25a, LC25a, Ran25]. value-at-risk [CLY25a]. valued [HY25a, MG25]. values [JFP25]. vanishing [ZHF25]. Variable [LS25, HYZ^{+}25, MS25, Sai25]. Variable-stepsize [LS25]. variance [PW25]. variational [Buo25, LY25, OAO25, TSM25, XHLH25]. variational-hemivariational [XHLH25]. varying [KPTS25, LSCZ25, SC25, XDZ25]. vector [TSM25]. vertical [FWL25, HV25]. via [BRPP25, CGR25, GBB25, KSC25, KIF^{+}25, QWLF25, SZD25, Yu25, ZC25b]. virtual [CG25, WCW25]. viscoelastic [LWL^{+}25]. Viscosity [LZZ25a]. Viscosity-independent [LZZ25a]. volatility [HKY25, Zho25]. volatility-barrier [HKY25]. Volterra [AFA25, AD25, FADA25, GHD25, MS25, RM25]. volume [LFH25, PY25].

wave [JS25, LWL^{+}25, Sak25]. wavelets [SKM25]. Weak [SHL^{+}25, ADJL25, LMZ25, QWZ25, Ton25]. weakly [KIF^{+}25, MS25]. weighted [BRPP25]. Well [XHLH25, AK25a]. well-known [AK25a]. Well-posedness [XHLH25]. wheeled [MZL25]. white [CTZ25, Ran25]. wise [MRRÁY25]. without [JJM25]. WKB [AKKM25]. Wolfe [ZHLD25].

XL [Kar25a].

Yosida [RNC25].

zeroing [MZL25]. zeros [FA25].

References

Alsenafi:2025:GAP

[AAN25] Abdulaziz Alsenafi, Fares Alazemi, and Alireza Najafi. Geometric Asian power option pricing with transaction cost under the geometric fractional Brownian motion with w sources of risk in fuzzy environment. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400414X.

Alghamdi:2025:GMM

[ABB25] Moataz Alghamdi, Daniele Boffi, and Francesca Bonizzoni. A greedy MOR method for the tracking of eigensolutions to parametric elliptic PDEs. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005193.

Alikhanov:2025:MMR

[ABH^{+}25] Anatoly Alikhanov, Huiran Bai, Jian Huang, Aleksei Tyrylgin, and Yin Yang. Multiscale model reduction for the time fractional thermoporoelasticity problem in fractured and heterogeneous media. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004060.

Arias:2025:HQH

[ACO25] Luis A. Souto Arias, Pasquale Cirillo, and Cornelis W. Oosterlee. The Heston–Queue– Hawkes process: a new selfexciting jump-diffusion model for options pricing, and an extension of the COS method for discrete distributions. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004266.

Aourir:2025:NCT

[AD25] E. Aourir and H. Laeli Dastjerdi. Numerical computational technique for solving Volterra integro-differential equations of the third kind using meshless collocation method. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005429.

Abreu:2025:SDL

[ADJL25] Eduardo Abreu, Richard De la cruz, Juan Juajibioy, and Wanderson Lambert. Semidiscrete Lagrangian–Eulerian approach based on the weak asymptotic method for nonlocal conservation laws in several dimensions. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005739.

Ahmad:2025:FPA

[ADY25] Haroon Ahmad, Fahim Ud Din, and Mudasir Younis. A fixed point analysis of fractional dynamics of heat transfer in chaotic fluid layers. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003935.

Amirali:2025:SON

[AFA25] Ilhame Amirali, Burcu Fedakar, and Gabil M. Amiraliyev. Second-order numerical method for a neutral Volterra integrodifferential equation. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004096.

Astorino:2025:DCP

[AGGM25] Annabella Astorino, Manlio Gaudioso, Enrico Gorgone, and Benedetto Manca. Difference of convex programming in adversarial SVM. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004503.

Abedi:2025:ESO

[AK25a] Hadis Abedi and Behrouz Kheirfam. An efficient secondorder predictor-corrector infeasible primal-dual IPM algorithm with large iteration path updates for solving wellknown SDO problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006277.

Ahmad:2025:RPE

[AK25b] Sk. Safique Ahmad and Pinki Khatun. A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400606X.

Arnold:2025:OTW

[AKKM25] Anton Arnold, Christian Klein, Jannis Körner, and Jens Markus Melenk. Optimally truncated WKB approximation for the 1D stationary Schrödinger equation in the highly oscillatory regime. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004898.

Arnal:2025:BSS

[AM25] A. Arnal and J. Monterde. Bézier-Smart surfaces of arbitrary degree. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005028.

Almutairi:2025:NEE

[AMRS25] Hanan Almutairi, Gérard Meurant, Lothar Reichel, and Miodrag M. Spalević. New error estimates for the conjugate gradient method. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006058.

Anonymous:2025:Ja

[Ano25a] Anonymous. 1 January 2025.

Journal of Computational and Applied Mathematics, 453(??): ??, January 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Ma

[Ano25b] Anonymous. 1 March 2025.

Journal of Computational and Applied Mathematics, 456(??): ??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Md

[Ano25c] Anonymous. 1 May 2025.

Journal of Computational and Applied Mathematics, 460(??): ??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jb

[Ano25d] Anonymous. 15 January 2025.

Journal of Computational and Applied Mathematics, 454(??): ??, January 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Mb

[Ano25e] Anonymous. 15 March 2025. Journal of Computational and Applied Mathematics, 457(??): ??, March 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Mc

[Ano25f] Anonymous. 15 May 2025. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Aa

[Ano25g] Anonymous. April 2025. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:EBa

[Ano25h] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S037704272400462X.

Anonymous:2025:EBb

[Ano25i] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004801.

Anonymous:2025:EBc

[Ano25j] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S037704272400493X.

Anonymous:2025:EBd

[Ano25k] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005156.

Anonymous:2025:EBe

[Ano25l] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005995.

Anonymous:2025:EBf

[Ano25m] Anonymous. Editorial Board.

Journal of Computational and Applied Mathematics, 458 (??):??, April 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042724006381.

Anonymous:2025:EBg

[Ano25n] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006824.

Anonymous:2025:EBh

[Ano25o] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042725000081.

Anonymous:2025:EBi

[Ano25p] Anonymous. Editorial Board. Journal of Computational and Applied Mathematics, 461 (??):??, June 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

tronic). URL http://www.
sciencedirect.com/science/
article/pii/S0377042725000512.

Anonymous:2025:F

[Ano25q] Anonymous. February 2025. Journal of Computational and Applied Mathematics, 455(??): ??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Anonymous:2025:Jc

[Ano25r] Anonymous. June 2025. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Aghamohammadi:2025:SRB

[AS25] Faezeh Aghamohammadi and Fatemeh Shakeri. Self representation based methods for tensor completion problem. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005454.

Baz:2025:GMR

[BAPPF25] Juan Baz, Pedro Alonso, Juan Manuel Peña, and Raúl Pérez-Fernández. Gaussian Markov Random Fields over graphs of paths and high relative accuracy. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003911.

Basak:2025:EAS

[BC25] Arpita Basak and Amit Choudhury. Estimating arrival and service rates from queue length observations. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400445X.

Boukraa:2025:FRM

[BCD25] Mohamed Aziz Boukraa, Laëtitia Caillé, and Franck Delvare. Fading regularization method for an inverse boundary value problem associated with the biharmonic equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005338.

Bauwelinckx:2025:CPC

[BDvdHV25] Yves-Cédric Bauwelinckx, Jan Dhaene, Milan van den Heuvel, and Tim Verdonck. On the causality-preservation capabilities of generative modelling. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005600.

Bazarra:2025:NAT

[BFQ25] N. Bazarra, J. R. Fernández, and R. Quintanilla. Numerical analysis of a thermoelastic problem of Moore–Gibson– Thompson type with history dependence in the thermal displacement. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400565X.

Bagheri:2025:MBH

[BKST25] Mahsa Bagheri, Faranges Kyanfar, Abbas Salemi, and Azita Tajaddini. A modified block Hessenberg method for low-rank tensor Sylvester equation. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004588.

Boroujeni:2025:NSK

[BPT25] Ahmad Aliyari Boroujeni, Reza Pourgholi, and Seyed Hashem Tabasi. Numerical solutions of KDV and mKDV equations: Using sequence and multicore parallelization implementation. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004333.

Butt:2025:NCJ

[BRPP25] S. I. Butt, T. Rasheed, D. Pecarić, and J. Pecarić. New cyclic Jensen type bounds via weighted HH inequalities and Montgomery identity with application to information inequalities. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003868.

Bandha:2025:NST

[BS25a] Raghunath Bandha and Rajen Kumar Sinha. A new space transformed finite element method for elliptic interface problems in R^{n}. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005260.

Bartel:2025:GOT

[BS25b] Andreas Bartel and Manuel Schaller. Goal-oriented time adaptivity for port-Hamiltonian systems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006988.

Bertolazzi:2025:PCR

[BS25c] Enrico Bertolazzi and Davide Stocco. Parallel cyclic reduction of padded bordered almost block diagonal matrices. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400579X.

Buong:2025:FOD

[Buo25] Nguyen Buong. A first order dynamical system and its discretization for a class of variational inequalities. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005892.

Chaskalovic:2025:EIA

[CA25a] Joël Chaskalovic and Franck Assous. Enhancing interpolation and approximation error estimates using a novel Taylorlike formula. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004023.

Chaskalovic:2025:RFO

[CA25b] Joël Chaskalovic and Franck Assous. A refined firstorder expansion formula in R^{n}: Application to interpolation and finite element error estimates. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005235.

Cai:2025:ASE

[CCLH25] Ting Cai, Pei Cheng, Xing Liu, and Mingang Hua. Almost sure exponential stability and stochastic stabilization of discrete-time stochastic systems with impulses. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004011.

Cuesta:2025:NAR

[CdlHG25] Carlota M. Cuesta, Francisco de la Hoz, and Ivan Girona. Numerical approximation of Riesz–Feller operators on R. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004436.

Chen:2025:UEA

[CG25] Yanping Chen and Jixiao Guo.

Unconditional error analysis of the linearized transformed L1 virtual element method for nonlinear coupled timefractional Schrödinger equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005272.

Cesarone:2025:ODM

[CGR25] Francesco Cesarone, Rosella Giacometti, and Jacopo Maria Ricci. Outlier detection of multivariate data via the maximization of the cumulant generating function. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724007052.

Chen:2025:TGF

[CH25] Yanping Chen and Hanzhang Hu. Two-grid finite element methods for space-fractional nonlinear Schrödinger equations. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006186.

Chen:2025:CMF

[Che25] Shuangshuang Chen. A combined mixed finite element method and discontinuous Galerkin method for hybriddimensional fracture models of two-phase flow. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006216.

Cai:2025:CIR

[CKC25] Xiao-Min Cai, Yi-Fen Ke, and Lin-Jie Chen. Color image restoration based on reduced biquaternion matrix singular value decomposition. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006769.

Chen:2025:USM

[CKL^{+}25] Meng Chen, Linghua Kong, Yiru Liang, Wei Wang, and Wei Yang. An unconditional stable modified finite element methods for Maxwell’s equation in Kerr-type nonlinear media. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004965.

Chao:2025:ERP

[CLJX25] Miantao Chao, Yongzi Lu, Jinbao Jian, and Xiao Xu. An efficient regularized PR splitting type algorithm for twoblock nonconvex linear constrained programs in l_{1/2} regularized compressed sensing problems. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724003947.

Chen:2025:NMP

[CLQY25] Jiawei Chen, Jie Liu, Xiaolong Qin, and Jen-Chih Yao. A non-monotone proximal gradient algorithm for solving nonsmooth multiobjective optimization problems with an extending application to robust multiobjective optimization. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006708.

Chen:2025:EUC

[CLY25a] Jie Chen, Haoxuan Li, and Xiangfeng Yang. An efficient uncertain chance constrained geometric programming model based on valueat-risk for truss structure optimization problems. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005958.

Chen:2025:ISN

[CLY25b] Pin-Bo Chen, Gui-Hua Lin, and Zhen-Ping Yang. An inexact semismooth Newton SAAbased algorithm for stochastic nonsmooth SOC complementarity problems with application to a stochastic power flow programming problem. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006095.

Chen:2025:CFP

[CLZ25] Ruyun Chen, Yu Li, and Yongxiong Zhou. On computation of finite-part integrals of highly oscillatory functions. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400582X.

Cano:2025:EER

[CM25] B. Cano and M. J. Moreta.

Efficient exponential Rosenbrock methods till order four. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004072.

Chen:2025:RHC

[CNQY25] An Chen, Thai Nguyen, Linyi Qian, and Zhixin Yang. The role of health in consumption and portfolio decision-making: Insights from state-dependent models. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005387.

Caballero:2025:CMF

[COPPS25] Josefa Caballero, Hanna Okrasińska-Płociniczak, Lukasz Płociniczak, and Kishin Sadarangani. Collocation method for a functional equation arising in behavioral sciences. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005910.

Chadjiconstantinidis:2025:BDM

[CP25]  Stathis  Chadjiconstantini-

dis and Konstadinos Politis. Bounds for the distribution and moments of the forward and backward recurrence times in a renewal process. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042724004151.

Cui:2025:GEM

[CQ25] Chunfeng Cui and Liqun Qi.

A genuine extension of the Moore–Penrose inverse to dual matrices. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004345.

Choi:2025:HPH

[CS25] Gary P. T. Choi and Mahmoud Shaqfa. Hemispheroidal parameterization and harmonic decomposition of simply connected open surfaces. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007039.

Cui:2025:FDM

[CTZ25] Yanzhen Cui, Shibing Tang, and Chao Zhang. Finite difference methods for stochastic Helmholtz equation driven by white noise. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S037704272400534X.

Cao:2025:DIH

[CWX25] Zhengbang Cao, Yimin Wei, and Pengpeng Xie. An L-DEIM induced high order tensor interpolatory decomposition. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724003923.

Cholamjiak:2025:DIS

[CXLP25] Prasit Cholamjiak, Zhongbing Xie, Min Li, and Papinwich Paimsang. Double inertial subgradient extragradient algorithm for solving equilibrium problems and common fixed point problems with application to image restoration. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006447.

Chen:2025:DPI

[CY25] Xuexue Chen and Cuilian You. Dynamic pricing and inventory model for perishable products with reference price and reference quality. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005867.

Du:2025:NES

[DdF25] Zhibin Du and Carlos M.

da Fonseca. A note on the eigenvalues of a Sylvester– Kac type matrix with offdiagonal biperiodic perturbations. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006770.

Diana:2025:NUS

[DG25] Derick Diana and Tim Gebbie. Non-uniformly sampled simulated price impact of an order-book. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004515.

Dembinska:2025:MLI

[DJ25] Anna Dembińska and Krzysztof Jasiński. Maximum likelihood inference about parameters of geometric lifetimes of heterogeneous components from data collected till failure of a k-out-of-n: G system. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004448.

Diegel:2025:ACD

[DLR25] Amanda E. Diegel, Xuejian Li, and Leo G. Rebholz. Analysis of continuous data assimilation with large (or even infinite) nudging parameters. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004709.

Frantsuzov:2025:GAB

[FA25] Viktor A. Frantsuzov and Anton V. Artemyev. A global argument-based algorithm for finding complex zeros and poles to investigate plasma kinetic instabilities. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004667.

Fedakar:2025:NSS

[FADA25] Burcu Fedakar, Ilhame Amirali, Muhammet Enes Durmaz, and Gabil M. Amiraliyev. Numerical solutions for second-order neutral Volterra integro-differential equations: Stability analysis and finite difference method. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006198.

Flavien:2025:JSP

[FBM25] Alonzo Flavien, Dia Ben Mansour, and Saad Mazen. Joint state-parameter estimation and inverse problems governed by reaction-advection-diffusion type PDEs with application to biological Keller– Segel equations and pattern formation. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007027.

Fahd:2025:EAP

[FBZ25] Hcini Fahd, Skander Belhaj, and Yulin Zhang. Efficient algorithms for perturbed symmetrical Toeplitz-plus-Hankel systems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005818.

Frasca-Caccia:2025:FDS

[FC25] Gianluca Frasca-Caccia. Finite difference schemes with non polynomial local conservation laws. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005788.

Fatone:2025:LCD

[FF25] Lorella Fatone and Daniele Funaro. Low-cost denoising and deblurring using a novel nonlinear diffusion technique. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400671X.

Faheem:2025:DDD

[FG25] Mo Faheem and Bapan Ghosh. Dynamics of a delayed discrete-time predator prey model proposed from a nonstandard finite difference scheme. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005946.

Ferreira:2025:NFD

[FGGdO25] José A. Ferreira, Mario Grassi, Elías Gudiño, and Paula de Oliveira. Non-Fickian diffusion enhanced by temperature. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005624.

Farouki:2025:AMC

[FKVZ25] Rida T. Farouki, Marjeta Knez, Vito Vitrih, and Emil Zagar. Application of a metric for complex polynomials to bounded modification of planar Pythagorean-hodograph curves. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004849.

Ferreira:2025:CAE

[FLS25] Chelo Ferreira, José L. López, and Ester Pérez Sinusía. Convergent and asymptotic expansions of the displacement elastodynamic integral in terms of known functions. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006435.

Fu:2025:NMB

[FWL25] Zu-Feng Fu, Shi-Liang Wu, and Liang Li. New modulusbased matrix splitting method for the vertical nonlinear complementarity problem. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005004.

Ghanem:2025:SDD

[GBB25] Aicha Ghanem, Touffik Bouremani, and Djamel Benterki. On the solution of Dolichobrachistochrone differential game via dynamic programming approach. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007088.

Gao:2025:NIR

[GCN25] Wansheng Gao, Ludovic Chamoin, and Insa Neuweiler. Non-intrusive reduced basis two-grid method for flow and transport problems in heterogeneous porous media. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005697.

Gidisu:2025:BDE

[GH25] Perfect Y. Gidisu and Michiel E.

Hochstenbach. Block discrete empirical interpolation methods. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004357.

Ghalini:2025:ESC

[GHD25] Rohollah Ghaedi Ghalini, Esmail Hesameddini, and Hojatollah Laeli Dastjerdi. An efficient spectral collocation method for solving Volterra delay integral equations of the third kind. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400387X.

Ge:2025:NEI

[GHZ25] Liang Ge, Jian Huang, and Jianwei Zhou. Novel efficient iterative schemes for linear finite element approximations of elliptic optimal control problems with integral constraint on the state. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004746.

Goroncy:2025:DTT

[GJ25] Agnieszka Goroncy and Krzysztof Jasiński. Discrete time threestate k-out-of-n system’s failure and numbers of components in each state. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005041.

Gong:2025:GBO

[GLTW25] Zhaohua Gong, Chongyang Liu, Kok Lay Teo, and Yonghong Wu. A gradientbased optimization algorithm to solve optimal control problems with conformable fractional-order derivatives. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004187.

Garcia:2025:NSS

[GNUV25] A. García, M. Negreanu, F. Ureña, and A. M. Vargas. On the numerical solution to space fractional differential equations using meshless finite differences. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005703.

Groselj:2025:RCP

[GP25] Jan Groselj and Ada Sadl Praprotnik. Rational C^{1} cubic Powell–Sabin B-splines with application to representation of ruled surfaces. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724005405.

Gu:2025:DCQ

[GR25] Yurun Gu and Clément Rey. Deterministic computation of quantiles in a Lipschitz framework. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005922.

Gudi:2025:EEB

[GS25] Thirupathi Gudi and Pratibha Shakya. Error estimates for a bubble-enriched C^{0} interior penalty method to a reduced state constrained elliptic optimal control problem. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006459.

Haghighi:2025:MSC

[Hag25]  Amir Haghighi.  Mean-

square convergence of an explicit derivative-free truncated method for nonlinear SDEs covering the non-commutative noise case. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006642.

Hou:2025:ISS

[HBG25] Liangliang Hou, Gongbing Bi, and Qianqian Guo. An improved sparrow search algorithm optimized LightGBM approach for credit risk prediction of SMEs in supply chain finance. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004461.

Honain:2025:GET

[HFSK25] Aljowhara H. Honain, Khaled M.

Furati, Ibrahim O. Sarumi, and Abdul Q. M. Khaliq. Generalized exponential time differencing for fractional oscillation models. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007040.

He:2025:IAA

[HHF25] Xin He, Rong Hu, and Ya-Ping Fang. Inertial accelerated augmented Lagrangian algorithms with scaling coefficients to solve exactly and inexactly linearly constrained convex optimization problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006733.

Han:2025:PMA

[HJ25] Wei-Wei Han and Yao-Lin Jiang. A penalty method for approximation of the stationary Stokes–Darcy problem. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005211.

Ham:2025:FOF

[HKH^{+}25] Seokjun Ham, Seungyoon Kang, Youngjin Hwang, Gyeonggyu Lee, Soobin Kwak, Jyoti, and Junseok Kim. A fourth-order finite difference method for the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004084.

Ha:2025:PTV

[HKY25] Mijin Ha, Donghyun Kim, and Ji-Hun Yoon. Pricing of timer volatility-barrier options under Heston’s stochastic volatility model. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005582.

Huang:2025:DAF

[HLL25] Yunqing Huang, Chanjie Li, and Jichun Li. Developing and analyzing a FDTD method for simulation of metasurfaces. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006125.

He:2025:PEI

[HLS25] Jun He, Yanmin Liu, and Xiaowei Shen. Pareto eigenvalue inclusion theorems and copositive tensors. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004424.

Howard:2025:CTI

[HRO^{+}25] Kyler Howard, Chris Rocheleau, Trevor Overton, Joel Barraza Nava, Mason Faldet, Kristina Moen, Summer Soller, Tyler Stephens, Esther van de Lagemaat, Natalie Wijesinghe, Kaylee Wong Dolloff, Nilton Barbosa da Rosa, and Jennifer L. Mueller. A comparison of techniques to improve pulmonary EIT image resolution using a database of simulated EIT images. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006630.

Huang:2025:SCS

[HRST25] Yu Huang, Narges Tohidi Rad, Mohammad Hadi Noori Skandari, and Emran Tohidi. A spectral collocation scheme for solving nonlinear delay distributed-order fractional equations. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400476X.

Hung:2025:SCC

[Hun25]  Nguyen Van Hung.  On

semi-continuity and continuity of solution maps of parametric generalized multiobjective games in fuzzy environments. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400551X.

Hussain:2025:GTF

[HuR25] Furqan Hussain and Mujeeb ur Rehman. On general tempered fractional calculus with Luchko kernels. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005879.

He:2025:NTM

[HV25] Jiewen He and Seakweng Vong. Newton-type methods for solving vertical linear complementarity problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006666.

He:2025:NPP

[HWZ25] Wenming He, Jiming Wu, and Zhimin Zhang. A novel post-processed finite element method and its convergence for partial differential equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005673.

Hu:2025:CIV

[HY25a] Xiang Hu and Jing Yao. A combined integer-valued autoregressive process with actuarial applications. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006320.

Hu:2025:SOL

[HY25b] Zuopeng Hu and Yanlong Yang. Stability of open-loop Nash equilibria for n-person nonzero-sum bounded rationality generalized differential games. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005090.

Huang:2025:NDD

[HYZ^{+}25] Jian Huang, Zhen Yue, Jing Zhang, Kai Fu, and Xilin Feng. Numerical discretization of a Darcy–Forchheimer flow with variable density and heat transfer. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005016.

Izuchukwu:2025:CIR

[IRS25] Chinedu Izuchukwu, Simeon Reich, and Yekini Shehu. Convergence of an inertial reflected-forward-backward splitting algorithm for solving monotone inclusion problems with application to image recovery. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006538.

Isik:2025:MMR

[ITM25] Osman Rasit Isik, Necibe Tuncer, and Maia Martcheva. A mathematical model for the role of vaccination and treatment in measles transmission in Turkey. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005569.

Jiang:2025:NAE

[JCCZ25] Ping Jiang, Jiejie Chen, Boshan Chen, and Zhigang Zeng. A novel adaptive eventtriggered controller for exponential quasi-synchronization and synchronization problems to coupled inertial memristive neural networks with time delays. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006745.

Jafari:2025:ALP

[JFP25] A. A. Jafari, S. Firoozi, and A. Pak. Analysis of the lifetime performance index in Gompertz distribution based on krecord values. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400400X.

Jian:2025:IBA

[JJM25] Jinbao Jian, Jiachen Jin, and Guodong Ma. An IDFPM-based algorithm without Lipschitz continuity to constrained nonlinear equations for sparse signal and blurred image restoration problems. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004837.

Jiang:2025:DGM

[JLC25] Xia Jiang, Rui Li, and Zhangxin Chen. A discontinuous Galerkin method for the Brinkman–Darcytransport problem. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004047.

Jornet:2025:PST

[JN25] Marc Jornet and Juan J. Nieto. The Peano–Sard theorem for fractional operators with Mittag-Leffler kernel and application in classical numerical approximation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724005119.

Janes:2025:NPO

[JS25] Andrew Janes and John R. Singler. A new proper orthogonal decomposition method with second difference quotients for the wave equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005284.

Jacob:2025:CQN

[JVW25] Birgit Jacob, Lukas Vorberg, and Christian Wyss. Computing the quadratic numerical range. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724002991.

Jiang:2025:THC

[JWS^{+}25] Meixuan Jiang, Yun Wang, Hu Shao, Ting Wu, and Weiwei Sun. Two hybrid conjugate gradient based algorithms on Riemannian manifolds with adaptive restart strategy for nonconvex optimization problems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007003.

Karageyik:2025:IFI

[Kar25a] Basak Bulut Karageyi̇k. An integrated framework for indicator-based decision analysis in proportional-XL reinsurance. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006897.

Kartal:2025:DFO

[Kar25b] Senol Kartal. A discrete fractional order Cournot duopoly game model with relative profit delegation: Stability, bifurcation, chaos, 0–1 testing and control. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005326.

Karageorghis:2025:MLM

[KC25] Andreas Karageorghis and C. S. Chen. Multi-level method of fundamental solutions for solving polyharmonic problems. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004692.

Kumar:2025:AFR

[KCM25] D. Kumar, A. K. B. Chand, and P. R. Massopust. Approximation with fractal radial basis functions. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004497.

Kattan:2025:SIE

[KIF^{+}25] Doha A. Kattan, Umar Ishtiaq, Muhammad Farhan, Ioannis K. Argyros, Mohammad Akram, and Samundra Regmi. Solution of an integral equation in controlled rectangular metric spaces via weakly contractive mappings. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004369.

Kang:2025:PPU

[KKHK25] Seungyoon Kang, Soobin Kwak, Youngjin Hwang, and Junseok Kim. Positivity preserving and unconditionally stable numerical scheme for the three-dimensional modified Fisher–Kolmogorov– Petrovsky–Piskunov equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005223.

Kalantarov:2025:BBD

[KKV25] V. K. Kalantarov, H. V.

Kalantarova, and O. Vantzos. Backward behavior and determining functionals for chevron pattern equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005314.

Kunwar:2025:DDL

[KL25] Hemanta Kunwar and Hyesuk Lee. Domain decomposition with local time discretization for the nonlinear Stokes– Biot system. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005594.

Kjelldahl:2025:NSG

[KM25] Vilma Kjelldahl and Ken Mattsson. Numerical simulation of the generalized modified Benjamin–Bona–Mahony equation using SBP-SAT in time. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006253.

Kaliraj:2025:FTS

[KPTS25] K. Kaliraj, P. K. Lakshmi Priya, V. Tamilarasan, and S. Suresh. Finite time stability of neutral multiterm fractional order time-varying delay systems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724007076.

Kiouvrekis:2025:DPS

[KS25a] Yiannis Kiouvrekis and Ioannis S. Stamatiou. Domain preserving and strongly converging explicit scheme for the stochastic SIS epidemic model. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004680.

Koellermeier:2025:PIM

[KS25b] Julian Koellermeier and Giovanni Samaey. Projective integration methods in the Runge–Kutta framework and the extension to adaptivity in time. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003960.

Kong:2025:FRR

[KS25c] Lingtao Kong and Chenwei Song. Fuzzy robust regression based on exponential-type kernel functions. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005430.

Katani:2025:ASM

[KSC25] Roghayeh Katani, Sedaghat Shahmorad, and Dajana Conte. Approximate solution of multi-term fractional differential equations via a block-by-block method. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003844.

Kouhanestani:2025:SIC

[KZG25] Mina Azizi Kouhanestani, Ehsan Zamanzade, and Sareh Goli. Statistical inference on the cumulative distribution function using judgment post stratification. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005880.

Li:2025:ENT

[LC25a] Xuehua Li and Cairong Chen. An efficient Newton-type matrix splitting algorithm for solving generalized absolute value equations with application to ridge regression problems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005776.

Lovat:2025:LDE

[LC25b] Giampiero Lovat and Salvatore Celozzi. Lucas decomposition and extrapolation methods for the evaluation of infinite integrals involving the product of three Bessel functions of arbitrary order. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400390X.

Li:2025:ACN

[LCZ25] Xiaolin Li, Xiyong Cui, and Shougui Zhang. Analysis of a Crank–Nicolson fast element-free Galerkin method for the nonlinear complex Ginzburg–Landau equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005715.

Lv:2025:KRM

[LF25] Xianli Lv and Xiufang Feng.

A kind regularization method for solving Cauchy problem of the Schrödinger equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004552.

Labarca-Figueroa:2025:CBV

[LFH25] Ignacio Labarca-Figueroa and Ralf Hiptmair. Coupled boundary and volume integral equations for electromagnetic scattering. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006915.

Li:2025:UBO

[LG25] Bo Li and Qinglong Gao. Uncertain bi-objective portfolio programming models of risky assets with liquidity and entropy constraints under uncertainty theory based DEA efficiency measures. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006903.

Linghu:2025:HOM

[LGDN25] Jiale Linghu, Weifeng Gao, Hao Dong, and Yufeng Nie. Higher-order multi-scale physics-informed neural network (HOMS-PINN) method and its convergence analysis for solving elastic problems of authentic composite materials. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004722.

Liu:2025:TSR

[LJY25] Wenli Liu, Jinbao Jian, and Jianghua Yin. A two-step relaxed-inertial derivative-free projection based algorithm for solving standard nonlinear pseudo-monotone equations and logistic regression problems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005752.

Liu:2025:MLE

[LLG25] Zhe Liu, Yang Liu, and Rong Gao. Maximum likelihood estimation for multi-factor uncertain differential equations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724004540.

Liu:2025:GRK

[LLXB25] Li Liu, Wei-Guo Li, Li-Li Xing, and Wen-Di Bao. Greedy Randomized Kaczmarz with momentum method for nonlinear equation. Journal of Computational and Applied Mathematics, 459(??): ??, May 15, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724006071.

Li:2025:SOL

[LLZH25] Yaoda Li, Min Li, Boya Zhou, and Zhibin Han. A second-order linear unconditionally energy-stable scheme for the phase field crystal equation. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004527.

Liu:2025:CAW

[LMZ25] Shicheng Liu, Xiangyun Meng, and Qilong Zhai. Convergence analysis of a weak Galerkin finite element method on a Shishkin mesh for a singularly perturbed fourth-order problem in 2D. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005727.

Lamwong:2025:FDM

[LP25a] Jiraporn Lamwong and Puntani Pongsumpun. A fractional derivative model of the dynamic of dengue transmission based on seasonal factors in Thailand. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005053.

Li:2025:CAA

[LP25b] Hongru Li and Miltiadis V.

Papalexandris. Convergence analysis and applicability of a domain decomposition method with nonlocal interface boundary conditions. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005259.

Loris:2025:CAP

[LR25] Ignace Loris and Simone Rebegoldi. Convergence analysis of a primal-dual optimization-by-continuation algorithm. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005478.

Li:2025:TGR

[LRG25] Hongpeng Li, Hongxing Rui, and Mengqi Gao. Two-grid reduced-order method based on POD for a nonlinear poroelasticity model. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006617.

Lang:2025:VSI

[LS25] Jens Lang and Bernhard A. Schmitt. Variable-stepsize implicit Peer triplets in ODE constrained optimal control. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006654.

Liu:2025:DIA

[LSCZ25] Xin Liu, Xianhua Song, Lili Chen, and Yanfeng Zhao. Distributed Ishikawa algorithms for seeking the fixed points of multi-agent global operators over time-varying communication graphs. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004990.

Los:2025:SIF

[LSPV25] Marcin Loś, Tomasz Słuzalec, Maciej Paszy’nski, and Eirik Valseth. Stabilization of isogeometric finite element method with optimal test functions computed from L_{2} norm residual minimization. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006587.

Li:2025:BEA

[LTZ25] Haibo Li, Guangming Tan, and Tong Zhao. Backward error analysis of the Lanczos bidiagonalization with reorthogonalization. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724006629.

Li:2025:ADG

[LW25] Xiaosheng Li and Wei Wang. An analysis of discontinuous Galerkin method for Electrical Impedance Tomography with partial data. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006241.

Liu:2025:RRT

[LWF25] Kai Liu, Bin Wang, and Ting Fu. Relaxation RKNtype integrators that preserve two invariants for second-order (oscillatory) systems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400548X.

Li:2025:CBS

[LWL^{+}25] Yaomeng Li, Feng Wang, Qiao Li, Chao Fu, and Xu Guo. Cubic B-spline based elastic and viscoelastic wave propagation method. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004850.

Lu:2025:DES

[LWZZ25] Nan Lu, Chenxi Wang, Lun Zhang, and Zhen Zhang. Decoupled and energy stable schemes for phase-field surfactant model based on mobility operator splitting technique. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006137.

Long:2025:SBG

[LY25] Xian-Jun Long and Jing Yang. A stochastic Bregman golden ratio algorithm for non-Lipschitz stochastic mixed variational inequalities with application to resource share problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006290.

Li:2025:ICE

[LYC25] Zishang Li, Changqing Ye, and Eric T. Chung. An iterative constraint energy minimizing generalized multiscale finite element method for contact problem. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006460.

Lancinskas:2025:PBA

[LZFP25] Algirdas Lancinskas, Julius Zilinskas, Pascual Fernández, and Blas Pelegrín. Populationbased algorithm for discrete facility location with ranking of candidate locations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005521.

Li:2025:VIR

[LZZ25a] Jingshi Li, Jiachuan Zhang, and Ran Zhang. Viscosityindependent reliability and efficiency constants for Brinkman equations with stabilized Crouzeix–Raviart element. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006599.

Lv:2025:RIA

[LZZ^{+}25b] Lingling Lv, Jiali Zhao, Bingqian Zheng, Jianwei Shen, and Huaicheng Yan. A recursive identification algorithm for discrete time-delay periodic linear systems. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006952.

Meng:2025:IAC

[MCL25] Xiaoqing Meng, Aijie Cheng, and Zhengguang Liu. Improving the accuracy and consistency of the energy quadratization method with an energyoptimized technique. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006162.

Mosic:2025:NEC

[MF25] D. Mosić and D. E. Ferreyra.

A new extension of the core inverse. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/pii/
S0377042724005533.

Mezzadri:2025:MBM

[MG25] Francesco Mezzadri and Emanuele Galligani. Modulus-based matrix splitting algorithms for generalized complex-valued horizontal linear complementarity problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006885.

Martin:2025:GPB

[MHO25] Dirk Martin, Gundolf Haase, and Günter Offner. Generic preconditioning based on the inverse with respect to the semi-inner product for iterative solvers for radial basis function interpolation. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006034.

Ma:2025:SMO

[MHWY25] Mengrao Ma, Linmin Hu, Lin Wang, and Xiaoyun Yu. Stochastic modeling and optimization of discrete-time cold standby repairable systems with unreliable repair facility and retrial mechanism. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003832.

Murad:2025:OSD

[MO25] Muhammad Amin S. Murad and Faraj M. Omar. Optical solitons, dynamics of bifurcation, and chaos in the generalized integrable (2 + 1)-dimensional nonlinear conformable Schrödinger equations using a new Kudryashov technique. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005466.

Moisa:2025:TOT

[Moi25] Andrew Moisa. Third order two-step Runge–Kutta– Chebyshev methods. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005399.

Masirevic:2025:ISR

[MPB25] Dragana Jankov Masirević, Tibor K. Pogány, and Tomislav Burić. Integral and series representations of the Nuttall function. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724003881.

Minakowski:2023:PPE

[MR23] P. Minakowski and T. Richter. A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem. Journal of Computational and Applied Mathematics, 421(??):??, March 15, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042722004435. See er-

ratum [MR25].

Minakowski:2025:EPP

[MR25] P. Minakowski and T. Richter. Erratum to “A priori and a posteriori error estimates for the Deep Ritz method applied to the Laplace and Stokes problem” [J. Comput. Appl. Math. 421 (2023) 114845]. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400654X.  See

[MR23].

Mahata:2025:GCN

[MRRÁY25] Shipra Mahata, Samala Rathan, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. A general correction for numerical integration rules over piece-wise continuous functions. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006265.

Ma:2025:NMB

[MS25] Zheng Ma and Martin Stynes.

A Nyström method based on product integration for weakly singular Volterra integral equations with variable exponent. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004138.

Masoodi:2025:OUS

[MT25] Mehdi Masoodi and Haniyeh Taromideh. Optimal and uniform sensor arrangement in near-field imaging. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004370.

Mao:2025:AOP

[MWTW25] Mengli Mao, Wansheng Wang, Tianhai Tian, and Lehan Wang. Adaptive option pricing based on a posteriori error estimates for fully discrete finite difference methods. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006551.

Miao:2025:NFT

[MZL25] Peng Miao, Daoyuan Zhang, and Shuai Li. A novel fixedtime zeroing neural network and its application to path tracking control of wheeled mobile robots. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006502.

Niu:2025:TAA

[NDSZ25] Jing Niu, Lei Du, Tomohiro Sogabe, and Shao-Liang Zhang. A tensor Alternating Anderson–Richardson method for solving multilinear systems with M-tensors. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006678.

Narbek:2025:CHP

[NKN25] Orazbek Narbek, Samat A.

Kassabek, and Targyn Nauryz. A collocation heat polynomials method for onedimensional inverse Stefan problems. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006046.

Nawa:2025:LMM

[NN25] Victor Nawa and Saralees Nadarajah. Logarithmic method of moments estimators for the Fréchet distribution. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005417.

Nikiforov:2025:MGM

[NPA^{+}25] Djulustan Nikiforov, Leonardo A. Poveda, Dmitry Ammosov, Yesy Sarmiento, and Juan Galvis. Meshfree generalized multiscale exponential integration method for parabolic problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006150.

Novo:2025:CNN

[NT25] Julia Novo and Eduardo Terrés. Can neural networks learn finite elements? Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724004175.

Niu:2025:RBK

[NZ25] Yu-Qi Niu and Bing Zheng. On the randomized block Kaczmarz algorithms for solving matrix equation AXB = C. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
S0377042724006691.

Oyewole:2025:ISE

[OAO25] Olawale K. Oyewole, Hammed A.

Abass, and Olufemi J. Ogunsola. An improved subgradient extragradient self-adaptive algorithm based on the golden ratio technique for variational inequality problems in Banach spaces. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400668X.

Orynyak:2025:TDE

[OTD^{+}25] Igor Orynyak, Anton Tsybulnyk, Kirill Danylenko, Andrii Oryniak, and Sergii Radchenko. Timestep-dependent element interpolation functions in the method of matched sections on the example of heat conduction problem. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004710.

Ozarslan:2025:AFC

[Öza25] Mehmet Ali Özarslan. On the approximation to fractional calculus operators with multivariate Mittag-Leffler function in the kernel. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003972.

Pecile:2025:DDD

[PDT^{+}25] Alessandro Pecile, Nicola Demo, Marco Tezzele, Gianluigi Rozza, and Dimitri Breda. Data-driven discovery of delay differential equations with discrete delays. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006873.

Phong:2025:RCH

[Pho25] Duong Thanh Phong. Representation computation for the hypergeometric function of a Hermitian matrix argument. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005077.

Park:2025:DGM

[PM25a] Do Yang Park and Minam Moon. Detective generalized multiscale hybridizable discontinuous Galerkin (GMsHDG) method for porous media. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005685.

Piterskaya:2025:SOS

[PM25b] Anna Piterskaya and Mikael Mortensen. A study of the Orr–Sommerfeld and induction equations by Galerkin and Petrov–Galerkin spectral methods utilizing Chebyshev polynomials. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006228.

Pelosi:2025:CPM

[PSF25] Francesca Pelosi, Maria Lucia Sampoli, and Rida T. Farouki. Control point modifications that preserve the Pythagorean-hodograph nature of planar quintic curves. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005491.

Peng:2025:ARM

[PW25] Xingchun Peng and Yankai Wang. Alpha-robust meanvariance reinsurance and investment strategies with transaction costs. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005065.

Peng:2025:ALE

[PY25] Gang Peng and Di Yang. An arbitrary Lagrangian– Eulerian positivity-preserving finite volume scheme for radiation hydrodynamics equations in the equilibrium-diffusion limit. Journal of Computational and Applied Mathematics, 455(??):??, February 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004059.

Qiu:2025:DME

[QWLF25] Wensong Qiu, Hongyan Wang, Yuan Li, and Lixin Feng. Determination of the modified exterior Steklov eigenvalues via the reciprocity gap method. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006083.

Qin:2025:IRP

[QWZ25] Mingze Qin, Xiuli Wang, and Huifang Zhou. Invariant region property of weak Galerkin method for semilinear parabolic equations. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006605.

Ranjbar:2025:EES

[Ran25]  Hassan Ranjbar.  An ex-

plicit Euler scheme for generalized n-dimensional secondorder differential equations with initial value conditions driven by additive Gaussian white noises. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400428X.

Roul:2025:HON

[RK25] Pradip Roul and Trishna Kumari. High-order numerical schemes based on B-spline for solving a time-fractional Fokker–Planck equation. Journal of Computational and Applied Mathematics, 460(??): ??, May 1, 2025. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S0377042724006344.

Ruby:2025:SRF

[RM25] Ruby and Moumita Mandal. Superconvergent results for fractional Volterra integrodifferential equations with non-smooth solutions. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005855.

Robert:2025:MYR

[RNC25] Maria Robert, Suresh Kumar Nadupuri, and Nagaiah Chamakuri. Moreau– Yosida regularization to optimal control of the monodomain model with pointwise control and state constraints in cardiac electrophysiology. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005545.

Saha:2025:IEP

[Sah25] Tanay Saha. Inverse eigenvalue problem for affinely structured or sparse polynomial matrices with prescribed eigenvalues. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004564.

Said:2025:IGR

[Sai25] Samia M. Said. Influence of gravitational on a rotating nonlocal thermoelastic medium with thermal variable conductivity. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004291.

Saksa:2025:CFE

[Sak25] Tytti Saksa. Comparison of finite element and discrete exterior calculus in computation of time-harmonic wave propagation with controllability. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004035.

Stoyanovskaya:2025:DAS

[SBAM25] O. P. Stoyanovskaya, O. A. Burmistrova, M. S. Arendarenko, and T. V. Markelova. Dispersion analysis of SPH for parabolic equations: Highorder kernels against tensile instability. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005648.

Shen:2025:RCL

[SC25] Zixiang Shen and Guo Chen. Robust H_{∞} control for LFC of discrete T–S fuzzy MAPS with DFIG and time-varying delays. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400520X.

Shekhar:2025:RAI

[SDS25] Chandra Shekhar, Mahendra Devanda, and Keshav Sharma. Reliability analysis of imperfect repair and switching failures: a Bayesian inference and Monte Carlo simulation approach. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007064.

Singh:2025:HEO

[SGRS25] Randhir Singh, Vandana Guleria, Higinio Ramos, and Mehakpreet Singh. Highly efficient optimal decomposition approach and its mathematical analysis for solving fourthorder Lane–Emden–Fowler equations. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004874.

Sugiura:2025:EBG

[SH25] Hiroshi Sugiura and Takemitsu Hasegawa. Error bounds for Gauss–Lobatto quadrature of analytic functions on an ellipse. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724005740.

Shadimetov:2025:OQF

[SHK25] Kholmat Shadimetov, Abdullo Hayotov, and Umedjon Khayriev. Optimal quadrature formulas for approximating strongly oscillating integrals in the Hilbert space W̃_{2}(m, m−1) of periodic functions. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724003820.

Shen:2025:WDB

[SHL^{+}25] Zhao-Li Shen, Guo-Liang Han, Yu-Tong Liu, Bruno Carpentieri, Chun Wen, and Jian-Jun Wang. Weak dangling block reordering and multistep block compression for efficiently computing and updating PageRank solutions. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005806.

Singh:2025:EWM

[SKM25] Akanksha Singh, Ankur Kanaujiya, and Jugal Mohapatra. Euler wavelets method for optimal control problems of fractional integro-differential equations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004278.

Shen:2025:LGM

[SNZ25] Binrui Shen, Qiang Niu, and Shengxin Zhu. Lightning graph matching. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004382.

Shang:2025:NPF

[SQLZ25] Youlin Shang, Deqiang Qu, Junxiang Li, and Roxin Zhang. A new parameterfree continuously differentiable filled function algorithm for solving nonlinear equations and data fitting problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004473.

Serna-Reyes:2025:CTS

[SRMGMD25] Adán J. Serna-Reyes, Siegfried Macías, Armando Gallegos, and Jorge E. Macías-Díaz. A convergent two-step method to solve a fractional extension of the Rosenau–Kawahara system. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006721.

Shen:2025:PTA

[SSLP25] Haoxuan Shen, Vladimir Simic, Shuguang Li, and Dragan Pamucar. Polynomial time algorithms to find Pareto optimal schedules of bicriteria lot scheduling problems with splitable jobs on a single parallel-batch machine. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006289.

Spiridonov:2025:PDO

[SSM25] Denis Spiridonov, Sergei Stepanov, and Tina Mai. Prediction of discretization of online GMsFEM using deep learning for Richards equation. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004163.

Sun:2025:GEB

[SWD25] Hongchun Sun, Yiju Wang, and Jiakang Du. Global error bound estimates algorithm for an R_{0}-type generalized LCP over polyhedral cone and its applications. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005363.

Song:2025:EAT

[SWWY25] Haiming Song, Hao Wang, Jiageng Wu, and Jinda Yang. An efficient ADAM-type algorithm with finite elements discretization technique for random elliptic optimal control problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004485.

Shah:2025:ASC

[SX25] Syed Omar Shah and Yonghui Xia. An analysis on the stability and controllability of non-linear measure differential equations. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724003893.

Song:2025:FFM

[SZD25] Qiu-Yan Song, Umair Zulfiqar, and Xin Du. Finitefrequency model order reduction of linear and bilinear systems via low-rank approximation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005351.

Tuncel:2025:SBS

[TA25] Altan Tuncel and Tugba Aktas Aslan. Some bivariate Schur-constant distributions and application to life insurance. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005442.

Tan:2025:SAE

[TC25] Ting Tan and Waixiang Cao. Spectral approximation and error analysis for the transmission eigenvalue problem with an isotropic inhomogeneous medium. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004126.

Tan:2025:ADB

[TGDC25] Longze Tan, Xueping Guo, Mingyu Deng, and Jingrun Chen. On the adaptive deterministic block Kaczmarz method with momentum for solving large-scale consistent linear systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005764.

Tong:2025:LSR

[Ton25] Hongzhi Tong. Least squares regression under weak moment conditions. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005843.

Treanta:2025:RSE

[TSM25] Savin Treanta, Vinay Singh, and Shashi Kant Mishra. Reciprocal solution existence results for a class of vector variational control inequalities with application in physics. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S037704272400709X.

Tschimpke:2025:RRD

[TST25] Marco Tschimpke, Manuela Schreyer, and Wolfgang Trutschnig. Revisiting the region determined by Spearman’s ρ and Spearman’s footrule φ. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042724005089.

Tan:2025:TSO

[TZ25a] Zhijun Tan and Yunhua Zeng. Temporal secondorder two-grid finite element method for semilinear time-fractional Rayleigh– Stokes equations. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400623X.

Tang:2025:LMT

[TZ25b] Jingyong Tang and Jinchuan Zhou. A Levenberg–Marquardt type algorithm with a Broydenlike update technique for solving nonlinear equations. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724006496.

Ullah:2025:CSH

[UX25] Inayat Ullah and Yingxiang Xu. Computational study of homogeneous–heterogeneous reactions in bioconvective third-grade material flow. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004308.

Vasilyeva:2025:GMF

[Vas25]  Maria Vasilyeva.  Gener-

alized multiscale finite element method for discrete network (graph) models. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005247.

Wang:2025:IMB

[WCS25] Lu-Xin Wang, Yang Cao, and Qin-Qin Shen. Improved modulus-based matrix splitting iteration methods for a class of horizontal implicit complementarity problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004321.

Wu:2025:EAT

[WCW25] Xiaohui Wu, Yanping Chen, and Yang Wang. Error analysis of two-grid virtual element method for nonlinear parabolic problems on general polygonal meshes. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006174.

Wang:2025:MFD

[WJ25] Ziyi Wang and Lijian Jiang. Model free data assimilation with Takens embedding. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006472.

Wang:2025:ELN

[WKQ25] Yingxiao Wang, Lingchen Kong, and Houduo Qi. An efficient Lagrange–Newton algorithm for long-only cardinality constrained portfolio selection on real data sets. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007015.

Wang:2025:CTS

[WLZZ25] Lixiu Wang, Huiyuan Li, Qian Zhang, and Zhimin Zhang. H(curl^{2})-conforming triangular spectral element method for quad-curl problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006101.

Wang:2025:CEH

[WX25] Weiwen Wang and Chuanju Xu. A class of efficient highorder time-stepping methods for the anisotropic phasefield dendritic crystal growth model. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004102.

Wang:2025:SKM

[WYZ25] Ze Wang, Jun-Feng Yin, and Ji-Chen Zhao. The sparse Kaczmarz method with surrogate hyperplane for the regularized basis pursuit problem. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400431X.

Wang:2025:ADC

[WZTL25] Liang Wang, Chunfang Zhang, Yogesh Mani Tripathi, and Yuhlong Lio. Analysis of dependent complementary competing risks data from a generalized inverted family of lifetime distributions under a maximum ranked set sampling procedure with unequal samples. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005570.

Xiao:2025:FOE

[XC25] Ming Xiao and Rui Chen.

A first-order energy stable scheme for the Allen–Cahn equation with the Allen–Cahn type dynamic boundary condition. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006575.

Xiong:2025:NFS

[XDZ25] Lianglin Xiong, Junzhou Dai, and Haiyang Zhang. Nonfragile sampled-data control for uncertain fractional-order systems with time-varying delay. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006861.

Xu:2025:NMF

[XGX25] Fei Xu, Yasai Guo, and Manting Xie. A novel multilevel finite element method for a generalized nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005296.

Xie:2025:TIG

[XGYH25] Wei Xie, Juan Galvis, Yin Yang, and Yunqing Huang. On time integrators for Generalized Multiscale Finite Element Methods applied to advection– diffusion in high-contrast multiscale media. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006113.

Xu:2025:WPC

[XHLH25] Wei Xu, Weimin Han, Ting Li, and Ziping Huang. Wellposedness of a class of evolutionary variational-hemivariational inequalities in contact mechanics. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006149.

Xie:2025:SCE

[Xie25]  Pengcheng Xie.  Sufficient

conditions for error distance reduction in the l^{2}-norm trust region between minimizers of local nonconvex multivariate quadratic approximates. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003959.

Xiao:2025:DRB

[XJQ25] Zhi-Hua Xiao, Yao-Lin Jiang, and Zhen-Zhong Qi. Dimension reduction based on time-limited cross Gramians for bilinear systems. Journal of Computational and Applied Mathematics, 457 (??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724005508.

Xu:2025:CSP

[XLLJ25] Yuncheng Xu, Sanyang Liu, Lixia Liu, and Kewei Jie. CVaR stochastic programming model for monotone stochastic tensor complementarity problem by using its penalized sample average approximation algorithm. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0377042724006757.

Xu:2025:DCN

[XLZ25] Bo Xu, Xinliang Liu, and Lei Zhang. Dilated convolution neural operator for multiscale partial differential equations. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006563.

Xin:2025:ASH

[XNL25] Xiaofang Xin, Guyan Ni, and Ying Li. An alternating shifted higher order power method based algorithm for rank-R Hermitian approximation and solving Hermitian CP-decomposition problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006332.

Xu:2025:UMC

[XQW25] Min Xu, Zhongfeng Qin, and Junbin Wang. Uncertain cmeans clustering method with application to imprecise observations. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005934.

Yao:2025:MNM

[Yao25]  Xudong Yao.  An MP-

Newton method for computing nonlinear eigenpairs and its application for solving a semilinear Schrödinger equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005636.

Yu:2025:CAR

[YC25] Dansheng Yu and Feilong Cao. Construction and approximation rate for feedforward neural network operators with sigmoidal functions. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003996.

Yin:2025:CSC

[Yin25]  Ruijie Yin.  Convergence

of some conservative Strang splitting methods for chargedparticle dynamics under a strong magnetic field. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006782.

Yousef:2025:PBL

[You25] Soleiman Yousef. A posterioribased, local multilevel mesh refinement for the Darcy porous media flow problem. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004114.

Yu:2025:PNR

[YPLH25] Xiang Yu, Yehui Peng, Penglin Lou, and Bozhong Huang. Poisson noise removal based on non-convex hybrid regularizers. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005375.

Yan:2025:RAD

[YSP25] Zaizai Yan, Yanjie Shi, and Xiuyun Peng. Reliability analysis of k/n(G) degradation system under dependent competing failures. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006927.

Yang:2025:PEE

[YTT25] Shuo Yang, Liutao Tian, and Hongjiong Tian. A posteriori error estimates and adaptivity for the IMEX BDF2 method for nonlinear parabolic equations. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005661.

Yu:2025:SNN

[Yu25] Yuexin Yu. Solving nonlinear neutral delay integrodifferential equations via general linear methods. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005909.

Yao:2025:EAM

[YWW25] Guoqing Yao, Zicheng Wang, and Zhongqing Wang. An efficient and accurate mapping method for elliptic equations in irregular annular domains. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004862.

Yang:2025:TMT

[YWY25] Yang Yang, Guojing Wang, and Jing Yao. Tail moments and tail joint moments for multivariate generalized hyperbolic distribution. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724005557.

Yan:2025:BIN

[YXM25] Xiong-Bin Yan, Zhi-Qin John Xu, and Zheng Ma. Bayesian Inversion with Neural Operator (BINO) for modeling subdiffusion: Forward and inverse problems. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004400.

Yao:2025:NCS

[YY25] Jinran Yao and Zhengwei Yin. Numerical contractivity of split-step backward Milstein-type schemes for commutative SDEs with nonglobally Lipschitz continuous coefficients. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006976.

Zhou:2025:NQG

[ZBK25] Jun Zhou, Pieter J. Barendrecht, and Jirí Kosinka. Numerical quadrature for Gregory triangles. Journal of Computational and Applied Mathematics, 453 (??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724003984.

Zhang:2025:MFE

[ZBN^{+}25] Zhizhuo Zhang, Mikaël Barboteu, Xiaobing Nie, Serge Dumont, Mahmoud Abdel-Aty, and Jinde Cao. Mixed finite element method for multi-layer elastic contact systems. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005302.

Zhang:2025:OEE

[ZC25a] Zhuoyue Zhang and Wentao Cai. Optimal L^{2} error estimates of mass- and energyconserved FE schemes for a nonlinear Schrödinger-type system. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005612.

Zhou:2025:QHO

[ZC25b] Yongxiong Zhou and Ruyun Chen. On quadrature of highly oscillatory Bessel function via asymptotic analysis of simplex integrals. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004886.

Zhukova:2025:ANA

[ZFCM25] Aleksandra Zhukova, Anna Flerova, Alexey Chernov, and Georgy Milyuchikhin. Approaches to numerical analysis of optimal control with linear phase constraints on the example of the assets and liabilities management by a bank. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724003790.

Zhang:2025:IIN

[ZH25] Xinru Zhang and Qiaolin He.

An immersed interface neural network for elliptic interface problems. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006204.

Zhu:2025:FCR

[ZHF25] Ting Ting Zhu, Rong Hu, and Ya Ping Fang. Fast convergence rates and trajectory convergence of a Tikhonov regularized inertial primal-dual dynamical system with time scaling and vanishing damping. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006423.

Zhu:2025:EMC

[ZHLD25] Zhibin Zhu, Jiaqi Huang, Ying Liu, and Yuehong Ding. An efficient modified conjugate gradient algorithm under Wolfe conditions with applications in compressive sensing. Journal of Computational and Applied Mathematics, 458(??):??, April 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005831.

Zhou:2025:PEO

[Zho25] Ke Zhou. Pricing exchange options under hybrid stochastic volatility and interest rate models. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724005107.

Zhang:2025:EFO

[ZLC^{+}25] Yanzhu Zhang, Tingting Liu, Yangquan Chen, Jing Wang, and Mingyu Shi. An efficient fractional-order PDE based image denoising algorithm with optimal adaptive strategy for ultrasound medical imagebased diagnostics. Journal of Computational and Applied Mathematics, 460(??):??, May 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006484.

Zhang:2025:DIE

[ZLS25] Zecheng Zhang, Wing Tat Leung, and Hayden Schaeffer. A discretization-invariant extension and analysis of some deep operator networks. Journal of Computational and Applied Mathematics, 456(??):??, March 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004758.

Zeng:2025:IMU

[ZLY25] Min Zeng, Xianlu Liao, and Yongxin Yuan. An iterative method for updating undamped structural systems with connectivity constraints. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004394.

Zhang:2025:ADS

[ZLZ25] Xuexue Zhang, Sanyang Liu, and Nannan Zhao. An accelerated decentralized stochastic optimization algorithm with inexact model. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724006319.

Zhang:2025:ECS

[ZMG25] Xiaohao Zhang, Liquan Mei, and Shimin Guo. Energyconserving SAV–Hermite–Galerkin spectral scheme with time adaptive method for coupled nonlinear Klein–Gordon system in multi-dimensional unbounded domains. Journal of Computational and Applied Mathematics, 454 (??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042724004539.

Zhu:2025:EMS

[ZNL25] Hong Zhu, Chenchen Niu, and Yongjin Liang. Efficient method for symmetric nonnegative matrix factorization with an approximate augmented Lagrangian scheme. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004679.

Zakharova:2025:CHA

[ZS25] Yulia Zakharova and Maria Sakhno. Complexity and heuristic algorithms for speed scaling scheduling of parallel jobs with energy constraint. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S037704272400503X.

Zhang:2025:EAE

[ZSYZ25] Jun Zhang, Fangying Song, Xiaofeng Yang, and Yu Zhang. Error analysis of the explicitinvariant energy quadratization (EIEQ) numerical scheme for solving the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 457(??):??, March 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724004734.

Zhou:2025:MIF

[ZWSC25] Sheng-Wei Zhou, Chun Wen, Zhao-Li Shen, and Bruno Carpentieri. The MFPIO iteration and the FPMPE method for multilinear PageRank computations. Journal of Computational and Applied Mathematics, 454(??):??, January 15, 2025. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724004412.

Zheng:2025:GIE

[ZZ25] Yawen Zheng and Song-Ping Zhu. A generalized integral equation formulation for pricing American options under regime-switching model. Journal of Computational and Applied Mathematics, 453(??):??, January 1, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724002668.

Zhang:2025:TSO

[ZZJJ25] Xin Zhang, Yiran Zhang, Qunzhi Jin, and Yuanfeng Jin. Two sixth-order, L_{∞} convergent, and stable compact difference schemes for the generalized Rosenau–KdV–RLW equation. Journal of Computational and Applied Mathematics, 459(??):??, May 15, 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0377042724006307.

Zhao:2025:TEM

[ZZX25] Jingjun Zhao, Hao Zhou, and Yang Xu. Two explicit methods for one-sided Lipschitz stochastic differential equations driven by fractional Brownian motion. Journal of Computational and Applied Mathematics, 461(??):??, June 2025. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0377042724007106.